### v1

# 4D Radar-Inertial Odometry based on Gaussian Modeling and Multi-Hypothesis Scan Matching

Abstract
4D millimeter-wave (mmWave) radars are sensors that provide robustness against adverse weather conditions (rain, snow, fog, etc.), and as such they are increasingly being used for odometry and SLAM applications. However, the noisy and sparse nature of the returned scan data proves to be a challenging obstacle for existing point cloud matching based solutions, especially those originally intended for more accurate sensors such as LiDAR. Inspired by visual odometry research around 3D Gaussian Splatting [8], in this paper we propose using freely positioned 3D Gaussians to create a summarized representation of a radar point cloud tolerant to sensor noise, and subsequently leverage its inherent probability distribution function for registration (similar to NDT [2]). Moreover, we propose simultaneously optimizing multiple scan matching hypotheses in order to further increase the robustness of the system against local optima of the function. Finally, we fuse our Gaussian modeling and scan matching algorithms into an EKF radar-inertial odometry system designed after current best practices. Experiments show that our Gaussian-based odometry is able to outperform current baselines on a well-known 4D radar dataset used for evaluation.
Our code and results can be publicly accessed at: https://github.com/robotics-upo/gaussian-rio
I Introduction
4D millimeter-wave (mmWave) radars have earned considerable popularity in the world of robotics as an alternative to classic camera or LiDAR sensors due to their robustness against adverse weather conditions (rain, snow, fog, etc.), as well as their compact size and low power consumption. As such, there is considerable interest in incorporating these sensors in odometry and SLAM applications. Existing LiDAR odometry solutions have been applied, adapted and sometimes even further specialized to 4D radar, with varying degrees of success [24, 25, 26, 22, 20]. However, the point clouds generated by 4D radars are sparser, noisier, and have a narrower field of view compared to those generated by LiDARs; and as a result these algorithms often struggle with the geometry of certain scenes. Therefore, it is desirable to explore scene representations that directly model coarse-grained surfaces and regions of radar point clouds instead of focusing on individual points themselves.
Another technique that has gained popularity recently in the field of robotics is 3D Gaussian Splatting (3DGS) [8]. This algorithm creates an explicit representation of a visual scene, and has been applied to SLAM [12, 7, 1] as an alternative to existing methods using implicit representations such as NeRF [13]. However, 3DGS heavily relies on the nature of visual information for training and rendering, and thus adapting it for use with LiDAR or radar instead of visual cameras is challenging; as opposed to NeRF, which can be straightforwardly adapted to model a SDF instead [14].
Figure 1: Gaussian model of an example point cloud, containing 150 Gaussians. The surface of each spheroid represents the $2\sigma$ front of each Gaussian (i.e. Mahalanobis distance equal to two).
Inspired by both of these lines of research, in this paper we propose and describe methods for obtaining a Gaussian model representation of a point cloud, and performing registration of a point cloud with respect to an existing Gaussian model. While this is similar in principle to the Normal Distributions Transform (NDT) [2], our methods consider Gaussians to be arbitrarily positioned in space instead of following a regular grid – like 3DGS [8]. Moreover, our scan matching (registration) step considers multiple simultaneous hypotheses, which can help increase the performance of the system, as we show in our experiments.
Figure 2: Overview of the proposed Gaussian radar-inertial odometry pipeline
Figure 2 shows the overall pipeline of our odometry system. As in other solutions, we use an Extended Kalman Filter (EKF) to perform inertial strapdown and process sensory input, including radar egovelocity. Keyframing is used to select individual 4D radar scans, which then undergo Gaussian modeling. The Gaussian multi-hypothesis scan matching algorithm registers every incoming 4D radar scan against the Gaussian model in order to produce a relative pose observation with respect to the last keyframe, which is then used to further update the EKF. Our experiments show how these Gaussian-based modeling and registration processes improve the results of existing state-of-the-art 4D radar-inertial odometry baselines.
This paper is organized as follows: Section II reviews existing related literature, Sections III, IV and V describe the three main components of our method respectively: Gaussian modeling, Gaussian scan matching, and the Extended Kalman Filter tying everything together; Section VI shows experimental results, and Section VII closes the paper with some conclusions.
II Related work
II-A Inertial odometry and SLAM
There exist a large number of works focused on Inertial Odometry, sometimes even attaining a full SLAM pipeline (including loop closure). Particularly we are interested in approaches that use non-visual sensors, such as [23, 21, 15, 3, 4, 24, 25, 26, 22, 20]. These works use a 3D LiDAR (LIO) or 4D radar (RIO) sensor together with an IMU to provide an odometry solution. The main technical advantage of these sensors over visual sensors (cameras) is their increased field/depth of view and robustness to adverse environmental conditions: in particular, both are immune to low light conditions, and 4D radar in addition offers robustness to fog, dust and rain. Moreover, 4D radar is uniquely capable of measuring the velocity of objects by leveraging the Doppler effect. Researchers have taken advantage of this to directly measure the robot’s own velocity (“egovelocity”) [3, 4], something which has been shown to significantly improve the performance of odometry systems.
Registration of incoming point clouds with respect to previous ones is a key technique used in odometry systems. In particular, the most widely used algorithm for aligning point clouds is Generalized Iterative Closest Point (GICP) [17]. In the case of radar scans, this is a more difficult task compared to LiDAR because of the lower number and density of points, as well as higher measurement noise/uncertainty in each point’s position. For this reason [26] proposed the Adaptive Probability Distribution variant of the algorithm (APDGICP), which leverages GICP’s ability to make use of the spatial probability distribution of each point, calculating the required point position covariance matrices according to the radar specifications. However, this is still in essence an approach focused on individual points, without attempting to extract larger scale geometric features that may be more stable and robust for registration purposes.
II-B Gaussian-based modeling
The Normal Distributions Transform (NDT) algorithm [2] slices and subdivides a point cloud using a regular grid and computes normal distributions for each cell that contains points. Scan matching is performed by maximizing the resulting probability distribution function with respect to an incoming scan. An important limitation of this algorithm is the fact that the grid structure itself imposes restrictions on the geometry, and the resulting model may not be fine grained enough to accurately model the geometry of the environment. Another problem lies in the fact that discontinuities in the probability function are produced when partially registered points cross cell boundaries during optimization. Moreover, the original authors only consider 2D scan matching (requiring 3D point clouds to be projected into the XY plane), although subsequent works [11] extend the algorithm to 3D. NDT has been used with radar scans too [10], although only for pure radar odometry as opposed to radar-inertial odometry, and without leveraging radar egovelocity information.
3D Gaussian Splatting (3DGS) [8] is a popular technique for creating a model of a scene from multiple posed images. A 3DGS model is a collection of multiple 3D normal distributions (Gaussians) freely distributed in space, each also containing color/radiance information encoded using spherical harmonics. Even though this technique was originally developed as an alternative to NeRF [13] for novel view synthesis, researchers have successfully leveraged this representation to develop SLAM solutions [12, 7, 1]. However, 3DGS and these works are all designed for visual sensors (cameras), and they cannot leverage other sensors such as LiDAR or radar due to the nature of the representation. This affects the rendering process in particular: while NeRF’s ray casting framework can be easily adapted to directly predict the output of a SDF [14] (therefore removing the volume rendering component from the system), the differentiable tile rasterizer is a core part of 3DGS that cannot be avoided, neither during training nor inference.
III Gaussian modeling
Given a point cloud $\mathcal{P}$ contaning $M$ points $\{\mathbf{p}_{i}|i=0,1,\ldots,M\}$ with $\mathbf{p}_{i}\in\mathbb{R}^{3}$, our goal is creating a summarized representation of its geometry for use with downstream tasks such as mapping or localization. Following the line of works such as NDT [2] and 3DGS [8], we decided to model the geometry using normal distributions. Figure 1 shows an example of a Gaussian model.
III-A Parametrization
We parametrize the Gaussian model $\mathcal{G}$ of a scene as a collection of $N$ independent trivariate normal distributions (3D Gaussians), each modeling the geometry of an individual region in space. The model is described by the following parameters of each Gaussian $\theta_{j}=\{\boldsymbol{\mu}_{j},\mathbf{s}_{j},\mathbf{q}_{j}\}$:
  • Center point $\boldsymbol{\mu}_{j}\in\mathbb{R}^{3}$, indicating the mean of the distribution.
  • Scaling vector $\mathbf{s}_{j}\in\mathbb{R}^{3}$, indicating the natural logarithms of the square roots of the eigenvalues of the covariance matrix.
  • Rotation quaternion $\mathbf{q}_{j}\in\mathbb{H}$, compactly encoding the eigenvectors of the covariance matrix (see below); where $\mathbb{H}$ is the field of quaternions.
Unlike in Gaussian Mixture Models (GMMs), we do not define $\pi_{j}$ weight parameters for the prior probability of a point being generated by a given Gaussian. This is because point cloud data does not encode a complete population of points – it merely represents a sample of said population, biased by factors such as the physical geometric properties of the sensor. For instance, sampling density is inversely proportional to distance from the sensor. This means we cannot model these prior probabilities, and as such we only consider posterior probabilities after a point is assigned to a Gaussian (through other non-probabilistic means, such as measuring distances).
The scaling and rotation parameters succinctly describe the covariance matrix of a normal distribution [19, 5]. To see why this is possible, we first derive a rotation matrix $\mathbf{R}_{j}=\mathbf{R}\{\mathbf{q}_{j}\}$, and a scaling matrix $\mathbf{S}_{j}=\mathrm{diag}(\exp_{\circ}(\mathbf{s}_{j}))$, where $\exp_{\circ}(\mathbf{x})$ is element-wise natural exponentiation and $\mathbf{R}\{\mathbf{q}\}$ is the quaternion-to-rotation-matrix conversion operator. We then define a transformation matrix $\mathbf{M}_{j}=\mathbf{R}_{j}\mathbf{S}_{j}$ that maps a mean-centered point into the Gaussian’s correlation space, and finally derive the covariance matrix $\boldsymbol{\Sigma}_{j}=\mathbf{M}_{j}\mathbf{M}_{j}^{T}$ of the distribution.
In addition, the auxiliary parameters $\hat{\mathbf{s}}_{j}$ and $\hat{\mathbf{q}}_{j}$ are defined and used during optimization, such that ${\hat{\mathbf{s}}_{j}=\max_{\circ}(s_{\mathrm{min}},\mathbf{s}_{j})}$ and $\mathbf{q}_{j}=\frac{\hat{\mathbf{q}}_{j}}{|\hat{\mathbf{q}}_{j}|}$ ($\max_{\circ}$ being the element-wise max operator). This transformation allows us to define a minimum log scale factor $s_{\mathrm{min}}$ to prevent Gaussians from becoming infinitely small (and Mahalanobis distances infinitely big); and also enforce that $\mathbf{q}_{i}$ must be a unit quaternion. Typically we make the minimum size equal to the radar point uncertainty (standard deviation of position noise).
III-B Initialization
The model $\mathcal{G}$ is initialized with all $\mathbf{s}_{j}$ set to 0 (indicating unit scale), and all $\hat{\mathbf{q}}_{j}$ set to the unit quaternion. As for $\boldsymbol{\mu}_{j}$, we perform a Bisecting K-Means clustering to find initial estimates for the centers of the Gaussians.
III-C Optimization
$\mathcal{G}$ is optimized iteratively using gradient descent. Each epoch involves the following steps:
  • Each point of $\mathcal{P}$ is matched to the nearest Gaussian center point $\boldsymbol{\mu}_{j}$. This process builds a set $G_{j}$ for each Gaussian, containing all matched points. We use the standard Euclidean distance in this step as opposed to the Mahalanobis distance in order to avoid a feedback loop between center and covariance matrix optimization that results in exploding gradients.
  • Each point $\mathbf{p}_{i}$ is transformed to the coordinate space of its assigned Gaussian as such: $\hat{\mathbf{p}}_{i}=\mathbf{M}_{j}^{-1}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})$. Note that $\mathbf{M}_{j}^{-1}=\mathbf{S}_{j}^{-1}\mathbf{R}_{j}^{T}$, and $\mathbf{S}_{j}^{-1}$ is trivial to calculate since $\mathbf{S}_{j}$ is a diagonal matrix.
  • The parameters of $\mathcal{G}$ (in other words, all $\theta_{j}$) are updated using the gradient of the loss function.
We define the following loss function affecting a given $\theta_{j}$:
$\mathcal{L}_{j}=\frac{1}{2|G_{j}|}\sum_{\mathbf{p}_{i}\in G_{j}}\hat{\mathbf{p}}_{i}^{T}\hat{\mathbf{p}}_{i}+\sum_{k}^{3}\hat{\mathbf{s}}_{j_{k}}+\left(\min_{k}^{3}\hat{s}_{j\,k}-s_{\textrm{disc}}\right)^{+},$ (1)
where $x^{+}=\max(0,x)$, and $s_{\textrm{disc}}$ is a hyperparameter that introduces a log size prior for the smallest dimension of the Gaussian, encouraging it to be shaped like a thin disc (better suited to modeling surfaces).
To see how the other terms of the loss function are derived, first we introduce the log PDF of the multivariate normal distribution, which we need to maximize, so we flip its sign:
$-\log\mathrm{P}(\mathbf{p}_{i}|\theta_{j})=K+\frac{1}{2}\log|\boldsymbol{\Sigma}_{j}|+\\ +\frac{1}{2}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})^{T}\boldsymbol{\Sigma}_{j}^{-1}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j}).$ (2)
[…]
$|\boldsymbol{\Sigma}_{j}|=|\mathbf{M}_{j}||\mathbf{M}_{j}^{T}|=|\mathbf{R}_{j}||\mathbf{S}_{j}|^{2}|\mathbf{R}_{j}^{T}|=|\mathbf{S}_{j}|^{2}=\left(\prod_{k}^{3}e^{\hat{s}_{j\,k}}\right)^{2},$ (3)
and thus,
$\frac{1}{2}\log|\boldsymbol{\Sigma}_{j}|=\frac{1}{2}2\sum_{k}^{3}\log e^{\hat{s}_{j\,k}}=\sum_{k}^{3}\hat{s}_{j\,k}.$ (4)
Likewise, we can simplify the term involving $p_{i}$. Note that it is half of the square of the Mahalanobis distance between the point and the Gaussian. The matrix $\boldsymbol{\Sigma}_{j}^{-1}$ can be decomposed as such: $\boldsymbol{\Sigma}_{j}^{-1}=(\mathbf{M}_{j}^{T})^{-1}\mathbf{M}_{j}^{-1}=(\mathbf{M}_{j}^{-1})^{T}\mathbf{M}_{j}^{-1}$. We can substitute it in the term and simplify:
$\frac{1}{2}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})^{T}\boldsymbol{\Sigma}_{j}^{-1}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})=\frac{1}{2}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})^{T}(\mathbf{M}_{j}^{-1})^{T}\cdot\\ \cdot\mathbf{M}_{j}^{-1}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})=\frac{1}{2}\hat{\mathbf{p}}_{i}^{T}\hat{\mathbf{p}}_{i}.$ (5)
Lastly, we convert the (per-point) log probability function into a per-Gaussian loss function by taking the mean log probability value across all points assigned to the Gaussian. This has the effect of introducing the $\frac{1}{|G_{j}|}$ factor in the Mahalanobis term, and has no effect in the determinant term since it is constant for all points within the Gaussian. We decided to aggregate with the mean in order to ensure that every Gaussian equally contributes towards the loss function, instead of being biased towards Gaussians with larger $G_{j}$ sets. The resulting loss function $\mathcal{L}$ of the entire model $\mathcal{G}$ can be finally obtained by once again taking the mean of the loss values across all Gaussians.
IV Gaussian multi-hypothesis scan matching
Given a Gaussian model $\mathcal{G}$ and a point cloud $\mathcal{P}$, we define a registration process that estimates the pose of the robot $\xi$ with respect to $\mathcal{G}$. The algorithm is based on the simultaneous optimization of multiple hypotheses centered around the current predicted position of the robot, in order to increase the robustness against local optima.
IV-A Definitions
We define the robot pose $\xi$ to be a member of $SE(3)$; that is, it includes a translational component $\mathbf{t}\in\mathbb{R}^{3}$ and rotational component $\mathbf{q}\in SO(3)$, with respect to some frame of reference. We consider that the uncertainty about the true robot pose follows some probability distribution: $\xi\thicksim\mathrm{P}$. Given a suitable model of $\mathrm{P}$, we can randomly sample a swarm of $K$ pose particles $\hat{\xi}^{k}$, each representing a different hypothesis, and use an optimization algorithm that evaluates each particle against $\mathcal{G}$ to further refine these estimates until one of the particles reaches a suitable optimum. We believe exploring the solution space using a particle swarm increases robustness in challenging situations, especially when dealing with large rotations or translations.
IV-B Optimization
The process is performed iteratively using gradient descent. Each epoch involves the following steps:
  • We define $K$ virtual copies of $\mathcal{P}$ called $\mathcal{P}^{\prime k}$, each registered according to every pose hypothesis $\hat{\xi}^{k}$. We also suggest downsampling the point cloud before creating the copies in order to reduce the computational load while retaining a representative sample of the geometry captured by the point cloud.
  • Every point $\mathbf{p}^{k}_{i}$ of $\mathcal{P}^{\prime k}$ is matched to the Gaussian $\theta_{j}$ in $\mathcal{G}$ that results in the lowest Mahalanobis distance, which we call $d^{k}_{i}$.
  • We calculate the loss function and perform gradient descent on $\hat{\xi}^{k}$.
The loss function used is simply
$\mathcal{L}_{k}=\frac{1}{M}\sum_{i}^{M}\min(d^{k}_{i},d_{\textrm{max}});\;\;\mathcal{L}=\sum_{i}^{K}\mathcal{L}_{i},$ (6)
where $d_{\textrm{max}}$ is the maximum Mahalanobis distance allowed, $\mathcal{L}_{k}$ is the loss function associated to a given particle $\hat{\xi}^{k}$, and $\mathcal{L}$ is the loss function affecting the entire swarm. Once the optimization process stops, we can select the particle with the lowest loss as the resulting output pose of the registration algorithm.
IV-C Keyframing
A common technique used in odometry systems is keyframing. Certain poses of the robot are selected to become local reference points (keyframes) that are later used for registration, graph optimization or loop closure purposes. In our case, we use keyframing to select individual radar point clouds to undergo Gaussian modeling, and serve as reference for Gaussian scan matching of subsequent radar point clouds. This has the benefit of reducing the accumulated error that arises from successive applications of the scan matching algorithm, though it also means the criteria used to select keyframe candidates needs to be carefully designed to generate the appropriate keyframe pacing. We use a simple set of criteria to decide when a new pose $\xi^{j}$ qualifies to become a new keyframe. Given the most recent keyframe $\xi^{i}$, we can derive the $SE(3)$ transformation relative to the keyframe $T=\xi^{j}\ominus\xi^{i}=\{\mathbf{t},\mathbf{q}\}$. The criteria are then simply that any of the following must hold:
  • For the translational component: $||\mathbf{t}||_{2}\geq d_{\max}$, where $d_{\max}$ is the maximum distance threshold.
  • For the rotational component: $2\cos^{-1}|q_{w}|\geq\alpha_{\max}$, where $\alpha_{\max}$ is the maximum rotation angle threshold.
V Inertial odometry with EKF
We use an Extended Kalman Filter in our odometry solution to tightly integrate all incoming sensor information: inertial data, radar egovelocity observations, and scan matching results. Unlike other published solutions [18, 3, 4], we only use Error State filtering for the attitude, as opposed to the entire state. The state ($\mathbf{x}$), control ($\mathbf{u}$) and noise ($\mathbf{w}$) vectors of the filter are the following:
$\mathbf{x}=\begin{bmatrix}\mathbf{p}\\ \mathbf{v}\\ \mathbf{b}_{a}\\ \mathbf{b}_{\omega}\\ \delta\boldsymbol{\theta}\end{bmatrix};\;\mathbf{u}=\begin{bmatrix}\mathbf{a}\\ \boldsymbol{\omega}\end{bmatrix};\;\mathbf{w}=\begin{bmatrix}\mathbf{w}_{v}\\ \mathbf{w}_{\theta}\\ \mathbf{w}_{a}\\ \mathbf{w}_{\omega}\\ \mathbf{w}_{b_{a}}\\ \mathbf{w}_{b_{\omega}}\end{bmatrix},$ (7)
where $\mathbf{p}$ is the position, $\mathbf{v}$ is the velocity, $\mathbf{b}_{a}$ and $\mathbf{b}_{\omega}$ are the accelerometer/gyroscope biases, $\delta\boldsymbol{\theta}$ is the attitude error expressed as a tangential rotation $\in\mathfrak{so}(3)$, and $\mathbf{a}$ and $\boldsymbol{\omega}$ are the IMU accelerometer/gyroscope readings. Velocity and attitude process noise are modeled by $\mathbf{w}_{v}$ and $\mathbf{w}_{\theta}$ respectively. Note that the attitude quaternion $\mathbf{q}$ is not included in $\mathbf{x}$, and the attitude error $\delta\boldsymbol{\theta}$ (which takes its place in $\mathbf{x}$) is nominally always 0. We follow Kalibr’s IMU noise model [16] [^1] https://github.com/ethz-asl/kalibr/wiki/IMU-Noise-Model, and consider both additive ($\mathbf{w}_{a}$, $\mathbf{w}_{\omega}$) and random walk noise ($\mathbf{w}_{b_{a}}$, $\mathbf{w}_{b_{\omega}}$) in accelerometer/gyroscope readings.
We employ the right-handed North-West-Up (NWU) convention typical of robotics, and consider three reference frames: world ($w$), body/IMU ($b$) and radar ($r$). The rotation matrix $\mathbf{C}^{g}_{f}$ transforms a vector from frame $f$ to frame $g$, likewise $\mathbf{C}^{f}_{g}=(\mathbf{C}^{g}_{f})^{T}$. We define $\mathbf{C}^{w}_{b}=\mathbf{R}\{\delta\mathbf{q}\}\mathbf{R}\{\mathbf{q}\}$, approximating $\mathbf{R}\{\delta\mathbf{q}\}=\exp\left([\delta\boldsymbol{\theta}]_{\times}\right)\approx\mathbf{I}+[\delta\boldsymbol{\theta}]_{\times}$ for the purposes of linearization, and $\mathbf{C}^{b}_{r}$ is given as a parameter. All state parameters ($\mathbf{p}$, $\mathbf{v}$, $\mathbf{q}$) are expressed in the world frame.
[…]
V-A EKF initialization
We initialize the state vector $\mathbf{x}$ to 0, attitude $\mathbf{q}$ to the unit quaternion, and covariance matrix $\mathbf{P}$ to 0 except for the following sub-blocks:
$\mathbf{P}_{b_{a}\,b_{a}}=\mathbf{I}\eta^{2}_{b_{a}};\;\mathbf{P}_{b_{\omega}\,b_{\omega}}=\mathbf{I}\eta^{2}_{b_{\omega}};\;\mathbf{P}_{\delta\theta_{xy}\,\delta\theta_{xy}}=\mathbf{I}\eta^{2}_{\theta},$ (8)
where $\eta_{b_{a}}$, $\eta_{b_{\omega}}$ and $\eta_{\theta}$ are the initial uncertainties (standard deviations) of the accelerometer bias, gyroscope bias and roll/pitch attitude error. We consider the initial yaw attitude uncertainty to be zero because we are not incorporating a compass/magnetometer.
V-B EKF propagation
We use a simplified inertial strapdown formulation that assumes movement to be linear instead of being perturbed by the current angular velocity reading of the gyroscope, due to the short time interval and small change in attitude (in effect equivalent to first order integration):
$\mathbf{p}$ $\leftarrow\mathbf{p}+\mathbf{v}t+\frac{1}{2}(\mathbf{C}^{w}_{b}(\mathbf{a}-\mathbf{b}_{a})+\mathbf{g})t^{2},$ (9)
$\mathbf{v}$ $\leftarrow\mathbf{v}+(\mathbf{C}^{w}_{b}(\mathbf{a}-\mathbf{b}_{a})+\mathbf{g})t,$ (10)
$\mathbf{q}$ $\leftarrow\mathbf{q}\otimes\exp\left(\frac{1}{2}(\boldsymbol{\omega}-\mathbf{b}_{\omega})t\right).$ (11)
The strapdown formulation is linearized in order to propagate the covariance matrix:
$\mathbf{P}\leftarrow\mathbf{F}\mathbf{P}\mathbf{F}^{T}+\mathbf{N}\mathbf{Q}\mathbf{N}^{T},$ (12)
where $\mathbf{F}$ is the linearized state transition matrix, $\mathbf{Q}$ is the noise covariance matrix, and $\mathbf{N}$ is the noise state transition matrix. These three matrices are sparsely populated: $\mathbf{F}=\mathbf{I}$ and $\mathbf{Q}=\mathbf{N}=\textbf{0}$, except for the following subblocks:
$\mathbf{F}_{p\,v}=\mathbf{I}t;\;\mathbf{F}_{p\,b_{a}}=-\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{F}_{p\,\delta\theta}=-\frac{1}{2}[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t^{2};$ (13)
$\mathbf{F}_{v\,b_{a}}=\mathbf{F}_{\delta\theta\,b_{\omega}}=-\mathbf{C}^{w}_{b}t;\;\mathbf{F}_{v\,\delta\theta}=-[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t;$ (14)
$\mathbf{Q}_{v\,v}=\mathbf{I}\sigma^{2}_{v};\;\mathbf{Q}_{\theta\,\theta}=\mathbf{I}\sigma^{2}_{\theta};\mathbf{Q}_{a\,a}=\mathbf{I}\frac{\sigma^{2}_{a}}{t};\;$ (15)
$\mathbf{Q}_{\omega\,\omega}=\mathbf{I}\frac{\sigma^{2}_{\omega}}{t};\mathbf{Q}_{b_{a}\,b_{a}}=\mathbf{I}\sigma^{2}_{b_{a}}t;\;\mathbf{Q}_{b_{\omega}\,b_{\omega}}=\mathbf{I}\sigma^{2}_{b_{\omega}}t;$ (16)
$\mathbf{N}_{p\,a}=\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{N}_{v\,a}=\mathbf{N}_{\delta\theta\,\omega}=\mathbf{C}^{w}_{b}t;$ (17)
$\mathbf{N}_{v\,v}=\mathbf{N}_{\delta\theta\,\theta}=\mathbf{N}_{b_{a}\,b_{a}}=\mathbf{N}_{b_{\omega}\,b_{\omega}}=\mathbf{I},$ (18)
where $\sigma_{v}$, $\sigma_{\theta}$, $\sigma_{a}$, $\sigma_{\omega}$, $\sigma_{b_{a}}$ and $\sigma_{b_{\omega}}$ are the standard deviations of the respective components of $\mathbf{w}$.
V-C EKF update and error state reset
Given observation $\mathbf{y}$, residual observation vector $\mathbf{r}$, observation covariance matrix $\mathbf{R}$ and linearized observation matrix $\mathbf{H}$ (so that $\mathbf{r}\approx\mathbf{y}-\mathbf{H}\mathbf{x}$), we can proceed to apply the Kalman filter equations (in Joseph form):
$\mathbf{K}=\mathbf{P}\mathbf{H}^{T}(\mathbf{H}\mathbf{P}\mathbf{H}^{T}+\mathbf{R})^{-1};\;\mathbf{L}=\mathbf{I}-\mathbf{K}\mathbf{H},$ (19)
$\mathbf{x}\leftarrow\mathbf{x}+\mathbf{K}\mathbf{r};\;\mathbf{P}\leftarrow\mathbf{L}\mathbf{P}\mathbf{L}^{T}+\mathbf{K}\mathbf{R}\mathbf{K}^{T}.$ (20)
As previously mentioned, the attitude part of the robot’s pose is expressed as error state in the EKF, and after an EKF update the nominal value of $\delta\boldsymbol{\theta}$ becomes non-zero, making it necessary to incorporate the estimated error back into the state as follows:
$\mathbf{q}\leftarrow\delta\mathbf{q}\otimes\mathbf{q};\;\delta\mathbf{q}=\exp\left(\frac{1}{2}\delta\boldsymbol{\theta}\right),$ (21)
$\delta\boldsymbol{\theta}\leftarrow 0;\;\mathbf{P}\leftarrow\mathbf{G}\mathbf{P}\mathbf{G}^{T};\;\mathbf{G}_{\delta\theta\,\delta\theta}=\mathbf{R}\{\delta\mathbf{q}\},$ (22)
where $\mathbf{G}$ is the error reset matrix, and it is a sparse matrix equal to $\mathbf{I}$ except for the given subblock.
V-D Radar egovelocity observation
Following previous works, we directly observe the robot’s velocity by leveraging the Doppler effect in measurements returned by the radar. The returned point cloud is first transformed by $\mathbf{C}_{r}^{b}$ to match the IMU orientation (we consider the separation between IMU and radar to be negligible). Afterwards, we use the RANSAC-LSQ algorithm [3] to both filter out outlier (dynamic) points, and estimate the current body-frame velocity of the robot, including a covariance matrix. The filtered point cloud is also fed to the Gaussian modeling and scan matching algorithm in the following subsection. The observation model used to perform the EKF update is as follows:
$\mathbf{r}=\mathbf{y}-\mathbf{C}^{b}_{w}\mathbf{v};\;\mathbf{H}_{:\,v}=\mathbf{C}^{b}_{w};\;\mathbf{H}_{:\,\delta\theta}=\mathbf{C}^{b}_{w}[\mathbf{v}]_{\times}.$ (23)
V-E Scan matching observation
The Gaussian modeling and scan matching algorithms allow us to observe robot poses relative to previous keyframes. In other words, given a keyframe pose $\xi^{k}$ we can observe the change in translation and rotation that has occurred since said keyframe. We initialize the pose particle swarm with a distribution $\mathbf{P}\thicksim\mathcal{N}(\xi^{x},\boldsymbol{\Sigma})$ where $\xi^{x}$ is the relative robot pose formed from the current nominal state of the EKF: $\xi^{x}=\{\mathbf{p},\mathbf{q}\}\ominus\xi^{k}$, and $\boldsymbol{\Sigma}$ is a covariance matrix indicating the desired dispersion of the particles for each of the 6 degrees of freedom (X/Y/Z/roll/pitch/yaw). Afterwards, we apply the scan matching algorithm and obtain an optimized observation pose $\xi^{y}$, from which a residual pose can be obtained: $\xi^{r}=\xi^{y}\ominus\xi^{x}=\{\delta\mathbf{p},\delta\mathbf{q}\}$. Since the attitude error quaternion $\delta\mathbf{q}$ can be approximated as $\delta\mathbf{q}=\exp\left(\frac{1}{2}\delta\boldsymbol{\theta}\right)\approx[1\,\frac{1}{2}\delta\boldsymbol{\theta}^{T}]^{T}$ for very small errors, the attitude error vector can be approximated as $\delta\boldsymbol{\theta}\approx\frac{2}{\delta q_{w}}\delta\mathbf{q}_{v}$. This allows us to construct the residual vector $\mathbf{r}=[\delta\mathbf{p}^{T}\,\delta\boldsymbol{\theta}^{T}]^{T}$ to be fed to the EKF update equation – note that this vector contains X/Y/Z/roll/pitch/yaw components (in this order). The non-zero subblocks of the observation matrix $\mathbf{H}$ can be defined as such: $\mathbf{H}_{\delta p\,p}=\mathbf{H}_{\delta\theta\,\delta\theta}=\mathbf{C}_{w}^{k}$, where $k$ corresponds to the keyframe’s reference frame. We are currently using a fixed observation covariance matrix $\mathbf{R}$, although we believe it should be possible to leverage the optimized particle swarm returned by the scan matching algorithm to estimate this matrix.
Furthermore, we constrain the observation model described above by removing degrees of freedom that are deemed noisy in the scan matching due to the properties of the sensor. For example, radars have lower resolution in the Z axis, which degrades the usefulness of Z/roll/pitch information returned by scan matching algorithms. In order to ignore the unreliable degrees of freedom, we transform the residual vector $\mathbf{H}_{c}\mathbf{r}$, observation matrix $\mathbf{H}_{c}\mathbf{H}$ and covariance matrix $\mathbf{H}_{c}\mathbf{R}\mathbf{H}_{c}^{T}$ by a constraint matrix defined as follows:
$\mathbf{H}_{c}=\begin{bmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&1\end{bmatrix}.$ (24)
Table I: Quantitative Analysis on NTU4DRadLM Dataset using EVO. Best results in bold, second best results underlined
cp nyl loop2 loop3
$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
(%) (º/m) (%) (º/m) (%) (º/m) (%) (º/m)
NTU4DRadLM baseline: gicp [25] 4.87 0.0648 5.65 0.0257 6.73 0.0094 6.03 0.0097
NTU4DRadLM baseline: Fast-LIO [25] 3.97 0.0773 4.47 0.0273 8.38 0.0096 6.95 0.0105
Our method (GICP scan matching) 1.69 0.0323 4.50 0.0198 3.92 0.0044 3.81 0.0054
Our method (single hypothesis) 1.47 0.0284 3.17 0.0150 2.95 0.0043 4.80 0.0065
Our method (full version) 1.33 0.0260 2.92 0.0134 3.49 0.0047 4.55 0.0061
Table II: Quantitative Analysis on NTU4DRadLM Dataset using rpg_trajectory_evaluation. Methods marked with * are self-reported metrics
cp nyl loop2 loop3
$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
(%) (º/m) (%) (º/m) (%) (º/m) (%) (º/m)
NTU4DRadLM baseline: gicp* [25] 4.13 0.0552 4.62 0.0184 4.84 0.0060 3.22 0.0060
NTU4DRadLM baseline: Fast-LIO* [25] 2.94 0.0468 3.80 0.0208 7.16 0.0057 4.55 0.0064
4DRadarSLAM apdgicp* [26] 3.56 0.0369 3.55 0.0171 6.09 0.0082 4.09 0.0097
EFEAR-4D* [22] 5.09 0.0125 8.93 0.0166 37.27 0.0152 37.01 0.0175
RIV-SLAM* [20] 2.58 0.0342 2.69 0.0456
NTU4DRadLM baseline: gicp [25] 4.921 0.068 5.775 0.027 6.743 0.010 6.242 0.010
NTU4DRadLM baseline: Fast-LIO [25] 4.178 0.083 4.759 0.030 8.221 0.010 7.120 0.011
Our method (GICP scan matching) 1.746 0.035 4.142 0.020 3.902 0.005 3.781 0.006
Our method (single hypothesis) 1.501 0.030 3.133 0.015 2.851 0.004 4.529 0.007
Our method (full version) 1.383 0.028 2.934 0.014 3.413 0.005 4.259 0.006
Figure 3: Generated trajectories for the 4 tested NTU4DRadLM sequences.
VI Experimental results
We compare our odometry with previous published 4D radar-inertial odometry methods on the NTU4DRadLM dataset [25]. These include the GICP and Fast-LIO baselines tested in the original NTU4DRadLM paper, the APDGICP algorithm tested in the follow-up 4DRadarSLAM paper [26], as well as other works such as EFEAR-4D [22] or RIV-SLAM [20]. We compare versions of these methods without loop closure if possible, in order to focus on evaluating pure odometry. Out of the six sequences present in NTU4DRadLM, we perform evaluation using four: cp and nyl (low speed regime, using a handcart platform), and loop2 and loop3 (high speed regime, using a car platform). The garden and loop1 sequences cannot be reliably used for evaluation because they contain large gaps in the IMU and radar data, respectively.
We evaluate relative translation/rotation errors using the modern and flexible evo evaluation package [6], as well as the older rpg_trajectory_evaluation package [27] used by previous works, using default settings. In the case of evo we use a custom evaluation script written specifically to follow the default behavior of rpg_trajectory_evaluation as closely as possible. We run both evaluation packages on the baseline GICP and Fast-LIO trajectories published along with the NTU4DRadLM dataset, as well as the trajectories generated by our odometry solution. In particular, we report results for the full version of our method, and also two ablated versions: one that replaces the Gaussian modeling and scan matching with GICP, and another that uses single-hypothesis scan matching. The GICP version uses small_gicp [9] as the underlying implementation, and in both ablations we use the current nominal state of the EKF directly as the initial estimate (instead of sampling a random particle swarm). We also include the self-reported metrics for all methods that appear in their respective papers as-is, since the trajectory files needed for metric calculation are not available.
The source code of our odometry solution can be accessed online on GitHub [^2] https://github.com/robotics-upo/gaussian-rio. We include the script used to generate the odometry trajectories, and the script used to evaluate trajectories with evo. The code also contains all details regarding the system parameters we used, such as EKF covariance initialization, IMU/noise parameters, number of Gaussians, etc.
VI-A Quantitative Analysis
VI-A1 Evaluation with evo
In Table I, we can observe major improvements in all 4 sequences provided by our method compared to the baselines originally tested by the authors of NTU4DRadLM. Moreover, the comparison between the Gaussian versions of our method and our GICP ablation shows improvements in relative translation error ($t_{rel}$) in 3 out of 4 sequences, and improvements in relative rotation error ($r_{rel}$) in 3 out of 4 sequences. In particular, the best improvements can be seen in the nyl sequence. The version with scan matching based on multiple hypotheses overperforms the version with a single hypothesis in 3 out of 4 sequences. We theorize that the sequences where our GICP ablation overperforms Gaussian versions of our method contain more challenging/sparser geometry due to being higher speed car driving sequences in open environments. Nonetheless, we believe it should be possible to further adjust the hyperparameters used (such as number of Gaussians used for modeling, number of hypotheses used for scan matching, keyframe thresholds, EKF covariances used, etc.) in order to optimize the algorithm for those sequences.
VI-A2 Evaluation with rpg_trajectory_evaluation
In Table II, we note discrepancies in the NTU4DRadLM baseline metrics between the reported values in the paper and the values obtained by running the evaluation software on the corresponding published trajectories, especially in the higher speed sequences. We also note very similar results for all methods generated by both evaluation packages, validating the use of both packages for evaluating 4D radar-inertial odometry methods. The errors reported by EFEAR-4D for the higher speed sequences are much higher than the values reported by other methods. We believe this to be caused by some difference in the evaluation methodology, as the same paper reports similar errors for baseline methods such as [26]. The most challenging sequence is loop3, and interestingly the best results are the ones reported by the NTU4DRadLM authors for the simplest baseline method. RIV-SLAM is the best performing method in loop2, however it must be said that to the best of our knowledge this method includes loop closure. The second best performing method in that sequence is the single-hypothesis version of our method, in line with the results in Table I.
VI-B Qualitative Analysis
Figure 3 shows the plotted trajectories of the ground truth, NTU4DRadLM GICP baseline and the full version of our method; across the 4 sequences. We can observe that in the low speed sequences (cp, nyl) our method accumulates substantially less drift in the XY plane, and manages to closely match the ground truth. High speed sequences (loop2, loop3) prove to be more challenging for both methods; in particular, our method drifts the most in loop3 with respect to the ground truth. On the other hand, our method performs much better than the baseline in the XZ plane across all 4 sequences. In particular, in loop2 and loop3 our method avoids large drifts in the Z axis, compared to the drifts in the baseline of over 70 meters. We think the higher accuracy in the XZ plane for our methods is responsible for improved quantitative results in these sequences, despite the lower accuracy in the XY plane.
VI-C Summary of results
We consider that the results described above validate our methodology. Our methods produce better relative translation and rotation metrics in 3 out of 4 tested sequences. Moreover, they produce generally better results regarding Z axis drift. Our ablation experiments also validate the Gaussian-based modeling and scan matching algorithm, as well as the multi-hypothesis approach; which contribute to improve both estimation accuracy and robustness.
VII Conclusion
In this paper we introduce a multi-hypothesis Gaussian-based radar-inertial odometry pipeline. Our methods summarize noisy 4D radar scans using freely positioned 3D Gaussians (like 3DGS [8] and unlike NDT [2]), which then support multi-hypothesis scan matching for increased robustness. We show that the fusion of these methods into an EKF produces better quantitative and qualitative results than existing baselines using traditional scan matching algorithms.
Future works include improving the Gaussian modeling and scan-matching algorithms (in particular, exploring optimization algorithms more direct than gradient descent), modeling kinematic restrictions relevant to the robot platform used in order to reduce noise during egovelocity estimation, further refining the pose estimates by means of Pose Graph Optimization (PGO), incorporating a loop closure module, and creating a global Gaussian map of the environment for the purposes of full SLAM.
Acknowledgment
This work was partially supported by the following grants: 1) INSERTION PID2021-127648OB-C31, and 2) NORDIC TED2021-132476B-I00 projects, funded by MCIN/AEI/ 10.13039/501100011033 and the “European Union NextGenerationEU / PRTR”.

### v2

@@
 Finally, we fuse our Gaussian modeling and scan matching algorithms into an EKF radar-inertial odometry system designed after current best practices.
-Experiments show that our Gaussian-based odometry is able to outperform current baselines on a well-known 4D radar dataset used for evaluation.
-Our code and results can be publicly accessed at: https://github.com/robotics-upo/gaussian-rio
+Experiments using publicly available 4D radar datasets show that our Gaussian-based odometry is comparable to existing registration algorithms, outperforming them in several sequences.
+Our code and results can be publicly accessed at: https://github.com/robotics-upo/gaussian-rio-cpp
 I Introduction
@@
 Figure 1: Gaussian model of an example point cloud, containing 150 Gaussians. The surface of each spheroid represents the $2\sigma$ front of each Gaussian (i.e. Mahalanobis distance equal to two).
-Inspired by both of these lines of research, in this paper we propose and describe methods for obtaining a Gaussian model representation of a point cloud, and performing registration of a point cloud with respect to an existing Gaussian model.
-While this is similar in principle to the Normal Distributions Transform (NDT) [2], our methods consider Gaussians to be arbitrarily positioned in space instead of following a regular grid – like 3DGS [8].
-Moreover, our scan matching (registration) step considers multiple simultaneous hypotheses, which can help increase the performance of the system, as we show in our experiments.
+In this paper we draw inspiration from both lines of research to propose the use of Gaussian models for representing a point cloud and performing registration with respect to it, as an alternative to existing registration methods such as the Normal Distributions Transform (NDT) [2] or Generalized Iterative Closest Point (GICP) [21].
+Gaussians are freely positioned in space instead of following a regular grid, as in 3DGS and unlike NDT.
+Moreover, our scan matching (registration) step considers multiple simultaneous hypotheses, which can in some situations increase the performance of the system, as we show in our experiments.
+In addition, we outline an Extended Kalman Filter (EKF) based approach to radar-inertial odometry that consolidates existing practices by integrating both radar egovelocity [4] and keyframing/scan matching components, which we then use to evaluate our Gaussian-based approach in comparison to other registration algorithms on multiple datasets and radar types.
 Figure 2: Overview of the proposed Gaussian radar-inertial odometry pipeline
@@
 The Gaussian multi-hypothesis scan matching algorithm registers every incoming 4D radar scan against the Gaussian model in order to produce a relative pose observation with respect to the last keyframe, which is then used to further update the EKF.
-Our experiments show how these Gaussian-based modeling and registration processes improve the results of existing state-of-the-art 4D radar-inertial odometry baselines.
 This paper is organized as follows: Section II reviews existing related literature, Sections III, IV and V describe the three main components of our method respectively: Gaussian modeling, Gaussian scan matching, and the Extended Kalman Filter tying everything together; Section VI shows experimental results, and Section VII closes the paper with some conclusions.
 II Related work
-II-A Inertial odometry and SLAM
-There exist a large number of works focused on Inertial Odometry, sometimes even attaining a full SLAM pipeline (including loop closure).
-Particularly we are interested in approaches that use non-visual sensors, such as [23, 21, 15, 3, 4, 24, 25, 26, 22, 20].
-These works use a 3D LiDAR (LIO) or 4D radar (RIO) sensor together with an IMU to provide an odometry solution.
+II-A Odometry and SLAM
+There exist a large number of works focused on odometry, sometimes even attaining a full SLAM pipeline (including loop closure).
+Particularly we are interested in approaches that use non-visual sensors, such as [27, 25, 19, 4, 5, 28, 29, 30, 26, 24, 31].
+These works use a 3D LiDAR or 4D radar sensor, sometimes together with an IMU, to provide an odometry solution.
 The main technical advantage of these sensors over visual sensors (cameras) is their increased field/depth of view and robustness to adverse environmental conditions: in particular, both are immune to low light conditions, and 4D radar in addition offers robustness to fog, dust and rain.
@@
 Researchers have taken advantage of this to directly measure the robot’s own velocity (“egovelocity”) [4, 5], something which has been shown to significantly improve the performance of odometry systems.
-Registration of incoming point clouds with respect to previous ones is a key technique used in odometry systems.
+Registration of incoming point clouds with respect to previous ones is a key technique used in some of the cited odometry systems.
 In particular, the most widely used algorithm for aligning point clouds is Generalized Iterative Closest Point (GICP) [21].
+This algorithm calculates distribution probability at each point based on their neighborhood, and optimizes a distribution-to-distribution scan matching function.
 In the case of radar scans, this is a more difficult task compared to LiDAR because of the lower number and density of points, as well as higher measurement noise/uncertainty in each point’s position.
@@
 However, this is still in essence an approach focused on individual points, without attempting to extract larger scale geometric features that may be more stable and robust for registration purposes.
+The works cited in this section use different sources of information, however many do not leverage all mentioned sources at the same time. Table I includes a list of surveyed methods.
 II-B Gaussian-based modeling
@@
 Moreover, the original authors only consider 2D scan matching (requiring 3D point clouds to be projected into the XY plane), although subsequent works [15] extend the algorithm to 3D.
-NDT has been used with radar scans too [10], although only for pure radar odometry as opposed to radar-inertial odometry, and without leveraging radar egovelocity information.
+NDT has been used with radar scans too [13, 14], albeit without inertial information.
+Another similar algorithm is the Voxelized GICP (VGICP) [12], which also creates an NDT-like voxel-based representation to serve as the reference model, and combines it with the distribution-to-distribution based approach of GICP.
 3D Gaussian Splatting (3DGS) [10] is a popular technique for creating a model of a scene from multiple posed images.
@@
 However, 3DGS and these works are all designed for visual sensors (cameras), and they cannot leverage other sensors such as LiDAR or radar due to the nature of the representation.
-This affects the rendering process in particular: while NeRF’s ray casting framework can be easily adapted to directly predict the output of a SDF [14] (therefore removing the volume rendering component from the system), the differentiable tile rasterizer is a core part of 3DGS that cannot be avoided, neither during training nor inference.
+This affects the rendering process in particular: while NeRF’s ray casting framework can be easily adapted to directly predict the output of a SDF [18] (therefore removing the volume rendering component from the system), the differentiable tile rasterizer is a core part of 3DGS that cannot be avoided, neither during training nor inference; thus requiring significant redesign efforts in order to synthesize other types of data such as LiDAR scans [3].
 III Gaussian modeling
@@
   • Scaling vector $\mathbf{s}_{j}\in\mathbb{R}^{3}$, indicating the natural logarithms of the square roots of the eigenvalues of the covariance matrix.
+We impose a minimum size restriction so that Gaussians do not become infinitely small (and Mahalanobis distances infinitely large).
   • Rotation quaternion $\mathbf{q}_{j}\in\mathbb{H}$, compactly encoding the eigenvectors of the covariance matrix (see below); where $\mathbb{H}$ is the field of quaternions.
+In order to ensure that the quaternions remain unitary during the optimization, we normalize them prior to usage.
 Unlike in Gaussian Mixture Models (GMMs), we do not define $\pi_{j}$ weight parameters for the prior probability of a point being generated by a given Gaussian.
@@
 We then define a transformation matrix $\mathbf{M}_{j}=\mathbf{R}_{j}\mathbf{S}_{j}$ that maps a mean-centered point into the Gaussian’s correlation space, and finally derive the covariance matrix $\boldsymbol{\Sigma}_{j}=\mathbf{M}_{j}\mathbf{M}_{j}^{T}$ of the distribution.
-In addition, the auxiliary parameters $\hat{\mathbf{s}}_{j}$ and $\hat{\mathbf{q}}_{j}$ are defined and used during optimization, such that ${\hat{\mathbf{s}}_{j}=\max_{\circ}(s_{\mathrm{min}},\mathbf{s}_{j})}$ and $\mathbf{q}_{j}=\frac{\hat{\mathbf{q}}_{j}}{|\hat{\mathbf{q}}_{j}|}$ ($\max_{\circ}$ being the element-wise max operator).
-This transformation allows us to define a minimum log scale factor $s_{\mathrm{min}}$ to prevent Gaussians from becoming infinitely small (and Mahalanobis distances infinitely big); and also enforce that $\mathbf{q}_{i}$ must be a unit quaternion.
-Typically we make the minimum size equal to the radar point uncertainty (standard deviation of position noise).
 III-B Initialization
-The model $\mathcal{G}$ is initialized with all $\mathbf{s}_{j}$ set to 0 (indicating unit scale), and all $\hat{\mathbf{q}}_{j}$ set to the unit quaternion.
-As for $\boldsymbol{\mu}_{j}$, we perform a Bisecting K-Means clustering to find initial estimates for the centers of the Gaussians.
+First, it is necessary to determine $N$, that is, the number of 3D Gaussians in the model.
+We propose choosing $N$ proportionally based on the number of points, so that richer scans can be represented with more Gaussians.
+The model $\mathcal{G}$ is then initialized with all $\mathbf{s}_{j}$ set to 0 (indicating unit scale), and all $\mathbf{q}_{j}$ set to the unit quaternion.
+As for $\boldsymbol{\mu}_{j}$, we perform Bisecting K-Means clustering to find initial estimates for the centers of the Gaussians.
 III-C Optimization
@@
   • The parameters of $\mathcal{G}$ (in other words, all $\theta_{j}$) are updated using the gradient of the loss function.
-We define the following loss function affecting a given $\theta_{j}$:
-$\mathcal{L}_{j}=\frac{1}{2|G_{j}|}\sum_{\mathbf{p}_{i}\in G_{j}}\hat{\mathbf{p}}_{i}^{T}\hat{\mathbf{p}}_{i}+\sum_{k}^{3}\hat{\mathbf{s}}_{j_{k}}+\left(\min_{k}^{3}\hat{s}_{j\,k}-s_{\textrm{disc}}\right)^{+},$ (1)
-where $x^{+}=\max(0,x)$, and $s_{\textrm{disc}}$ is a hyperparameter that introduces a log size prior for the smallest dimension of the Gaussian, encouraging it to be shaped like a thin disc (better suited to modeling surfaces).
-To see how the other terms of the loss function are derived, first we introduce the log PDF of the multivariate normal distribution, which we need to maximize, so we flip its sign:
-$-\log\mathrm{P}(\mathbf{p}_{i}|\theta_{j})=K+\frac{1}{2}\log|\boldsymbol{\Sigma}_{j}|+\\ +\frac{1}{2}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})^{T}\boldsymbol{\Sigma}_{j}^{-1}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j}).$ (2)
-[…]
-$|\boldsymbol{\Sigma}_{j}|=|\mathbf{M}_{j}||\mathbf{M}_{j}^{T}|=|\mathbf{R}_{j}||\mathbf{S}_{j}|^{2}|\mathbf{R}_{j}^{T}|=|\mathbf{S}_{j}|^{2}=\left(\prod_{k}^{3}e^{\hat{s}_{j\,k}}\right)^{2},$ (3)
-and thus,
-$\frac{1}{2}\log|\boldsymbol{\Sigma}_{j}|=\frac{1}{2}2\sum_{k}^{3}\log e^{\hat{s}_{j\,k}}=\sum_{k}^{3}\hat{s}_{j\,k}.$ (4)
-Likewise, we can simplify the term involving $p_{i}$.
-Note that it is half of the square of the Mahalanobis distance between the point and the Gaussian.
-The matrix $\boldsymbol{\Sigma}_{j}^{-1}$ can be decomposed as such: $\boldsymbol{\Sigma}_{j}^{-1}=(\mathbf{M}_{j}^{T})^{-1}\mathbf{M}_{j}^{-1}=(\mathbf{M}_{j}^{-1})^{T}\mathbf{M}_{j}^{-1}$.
-We can substitute it in the term and simplify:
-$\frac{1}{2}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})^{T}\boldsymbol{\Sigma}_{j}^{-1}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})=\frac{1}{2}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})^{T}(\mathbf{M}_{j}^{-1})^{T}\cdot\\ \cdot\mathbf{M}_{j}^{-1}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})=\frac{1}{2}\hat{\mathbf{p}}_{i}^{T}\hat{\mathbf{p}}_{i}.$ (5)
-Lastly, we convert the (per-point) log probability function into a per-Gaussian loss function by taking the mean log probability value across all points assigned to the Gaussian.
-This has the effect of introducing the $\frac{1}{|G_{j}|}$ factor in the Mahalanobis term, and has no effect in the determinant term since it is constant for all points within the Gaussian.
-We decided to aggregate with the mean in order to ensure that every Gaussian equally contributes towards the loss function, instead of being biased towards Gaussians with larger $G_{j}$ sets.
-The resulting loss function $\mathcal{L}$ of the entire model $\mathcal{G}$ can be finally obtained by once again taking the mean of the loss values across all Gaussians.
+We define the following loss function affecting a given $\theta_{j}$, derived from the log PDF of the multivariate normal distribution:
+$\mathcal{L}_{j}=\frac{1}{2|G_{j}|}\sum_{\mathbf{p}_{i}\in G_{j}}\hat{\mathbf{p}}_{i}^{T}\hat{\mathbf{p}}_{i}+\sum_{k}^{3}{\mathbf{s}_{j_{k}}.}$ (1)
+The loss function $\mathcal{L}$ of the entire model $\mathcal{G}$ can be obtained by taking the mean of the loss values across all Gaussians.
 IV Gaussian multi-hypothesis scan matching
@@
 IV-B Optimization
-The process is performed iteratively using gradient descent. Each epoch involves the following steps:
+The process is performed iteratively using Gauss-Newton. Each iteration involves the following steps:
   • We define $K$ virtual copies of $\mathcal{P}$ called $\mathcal{P}^{\prime k}$, each registered according to every pose hypothesis $\hat{\xi}^{k}$.
-We also suggest downsampling the point cloud before creating the copies in order to reduce the computational load while retaining a representative sample of the geometry captured by the point cloud.
   • Every point $\mathbf{p}^{k}_{i}$ of $\mathcal{P}^{\prime k}$ is matched to the Gaussian $\theta_{j}$ in $\mathcal{G}$ that results in the lowest Mahalanobis distance, which we call $d^{k}_{i}$.
-  • We calculate the loss function and perform gradient descent on $\hat{\xi}^{k}$.
-The loss function used is simply
-$\mathcal{L}_{k}=\frac{1}{M}\sum_{i}^{M}\min(d^{k}_{i},d_{\textrm{max}});\;\;\mathcal{L}=\sum_{i}^{K}\mathcal{L}_{i},$ (6)
-where $d_{\textrm{max}}$ is the maximum Mahalanobis distance allowed, $\mathcal{L}_{k}$ is the loss function associated to a given particle $\hat{\xi}^{k}$, and $\mathcal{L}$ is the loss function affecting the entire swarm.
-Once the optimization process stops, we can select the particle with the lowest loss as the resulting output pose of the registration algorithm.
+  • We perform the usual Gauss-Newton update step for each particle $\hat{\xi}^{k}$, using the sums of squared Mahalanobis distances as the function to optimize, as in GICP.
+We also exclude points whose $d^{k}_{i}$ is greater than a given distance threshold $d_{\textrm{max}}$.
+The optimization stops when all particles converge to a solution (i.e.
+when the update applied is under a given order of magnitude), or when a maximum number of iterations is reached; whichever happens first.
+The resulting output pose of the registration algorithm is the particle with the lowest value of the scoring function $\mathcal{L}_{k}$ defined as follows:
+$\mathcal{L}_{k}=\frac{1}{M}\sum_{i}^{M}\min(d^{k}_{i},d_{\textrm{max}}){,}$ (2)
+where $M$ is the number of points in the scan. We consider a failure condition when the output particle has not converged.
 IV-C Keyframing
@@
   • For the rotational component: $2\cos^{-1}|q_{w}|\geq\alpha_{\max}$, where $\alpha_{\max}$ is the maximum rotation angle threshold.
+As a fallback, if no scans have returned a successful match during the past given time interval $t$, a keyframe is automatically generated.
 V Inertial odometry with EKF
 We use an Extended Kalman Filter in our odometry solution to tightly integrate all incoming sensor information: inertial data, radar egovelocity observations, and scan matching results.
-Unlike other published solutions [18, 3, 4], we only use Error State filtering for the attitude, as opposed to the entire state.
-The state ($\mathbf{x}$), control ($\mathbf{u}$) and noise ($\mathbf{w}$) vectors of the filter are the following:
-$\mathbf{x}=\begin{bmatrix}\mathbf{p}\\ \mathbf{v}\\ \mathbf{b}_{a}\\ \mathbf{b}_{\omega}\\ \delta\boldsymbol{\theta}\end{bmatrix};\;\mathbf{u}=\begin{bmatrix}\mathbf{a}\\ \boldsymbol{\omega}\end{bmatrix};\;\mathbf{w}=\begin{bmatrix}\mathbf{w}_{v}\\ \mathbf{w}_{\theta}\\ \mathbf{w}_{a}\\ \mathbf{w}_{\omega}\\ \mathbf{w}_{b_{a}}\\ \mathbf{w}_{b_{\omega}}\end{bmatrix},$ (7)
-where $\mathbf{p}$ is the position, $\mathbf{v}$ is the velocity, $\mathbf{b}_{a}$ and $\mathbf{b}_{\omega}$ are the accelerometer/gyroscope biases, $\delta\boldsymbol{\theta}$ is the attitude error expressed as a tangential rotation $\in\mathfrak{so}(3)$, and $\mathbf{a}$ and $\boldsymbol{\omega}$ are the IMU accelerometer/gyroscope readings.
-Velocity and attitude process noise are modeled by $\mathbf{w}_{v}$ and $\mathbf{w}_{\theta}$ respectively.
-Note that the attitude quaternion $\mathbf{q}$ is not included in $\mathbf{x}$, and the attitude error $\delta\boldsymbol{\theta}$ (which takes its place in $\mathbf{x}$) is nominally always 0.
-We follow Kalibr’s IMU noise model [16] [^1] https://github.com/ethz-asl/kalibr/wiki/IMU-Noise-Model, and consider both additive ($\mathbf{w}_{a}$, $\mathbf{w}_{\omega}$) and random walk noise ($\mathbf{w}_{b_{a}}$, $\mathbf{w}_{b_{\omega}}$) in accelerometer/gyroscope readings.
+Unlike other published solutions [22, 4, 5], we only use Error State filtering for rotational variables, as opposed to the entire state.
 We employ the right-handed North-West-Up (NWU) convention typical of robotics, and consider three reference frames: world ($w$), body/IMU ($b$) and radar ($r$).
 The rotation matrix $\mathbf{C}^{g}_{f}$ transforms a vector from frame $f$ to frame $g$, likewise $\mathbf{C}^{f}_{g}=(\mathbf{C}^{g}_{f})^{T}$.
-We define $\mathbf{C}^{w}_{b}=\mathbf{R}\{\delta\mathbf{q}\}\mathbf{R}\{\mathbf{q}\}$, approximating $\mathbf{R}\{\delta\mathbf{q}\}=\exp\left([\delta\boldsymbol{\theta}]_{\times}\right)\approx\mathbf{I}+[\delta\boldsymbol{\theta}]_{\times}$ for the purposes of linearization, and $\mathbf{C}^{b}_{r}$ is given as a parameter.
-All state parameters ($\mathbf{p}$, $\mathbf{v}$, $\mathbf{q}$) are expressed in the world frame.
+We decompose each $\mathbf{C}$ into an error part and a nominal quaternion part as follows: $\mathbf{C}=\exp\left([\delta\boldsymbol{\theta}]_{\times}\right)\mathbf{R}\{\mathbf{q}\}$, approximating the error part as $\mathbf{I}+[\delta\boldsymbol{\theta}]_{\times}$ for the purposes of linearization.
+The state ($\mathbf{x}$), control ($\mathbf{u}$) and noise ($\mathbf{w}$) vectors of the filter are the following:
+$\mathbf{x}=\begin{bmatrix}\mathbf{p}\\ \mathbf{v}\\ {\mathbf{t}_{r}^{b}}\\ \mathbf{b}_{a}\\ \mathbf{b}_{\omega}\\ {\delta\boldsymbol{\theta}_{b}^{w}}\\ {\delta\boldsymbol{\theta}_{r}^{b}}\end{bmatrix};\;\mathbf{u}=\begin{bmatrix}\mathbf{a}\\ \boldsymbol{\omega}\end{bmatrix};\;\mathbf{w}=\begin{bmatrix}\mathbf{w}_{v}\\ \mathbf{w}_{\theta}\\ \mathbf{w}_{a}\\ \mathbf{w}_{\omega}\\ \mathbf{w}_{b_{a}}\\ \mathbf{w}_{b_{\omega}}\end{bmatrix},$ (3)
+where $\mathbf{p}$ is the position, $\mathbf{v}$ is the velocity, $\mathbf{t}_{r}^{b}$ is the translation between radar and IMU, $\mathbf{b}_{a}$ and $\mathbf{b}_{\omega}$ are the accelerometer/gyroscope biases, each $\delta\boldsymbol{\theta}$ is the error part of its corresponding $\mathbf{C}$ frame expressed as a tangential rotation $\in\mathfrak{so}(3)$, and $\mathbf{a}$ and $\boldsymbol{\omega}$ are the IMU accelerometer/gyroscope readings.
+The position and velocity of the robot are expressed in the world frame, while the nominal rotations $\mathbf{q}_{b}^{w}$ and $\mathbf{q}_{r}^{b}$ are stored outside the filter state as quaternions.
+Velocity and attitude process noise are modeled by $\mathbf{w}_{v}$ and $\mathbf{w}_{\theta}$ respectively.
+We follow Kalibr’s IMU noise model [20] [^1] https://github.com/ethz-asl/kalibr/wiki/IMU-Noise-Model, and consider both additive ($\mathbf{w}_{a}$, $\mathbf{w}_{\omega}$) and random walk noise ($\mathbf{w}_{b_{a}}$, $\mathbf{w}_{b_{\omega}}$) in accelerometer/gyroscope readings.
 […]
 V-A EKF initialization
-We initialize the state vector $\mathbf{x}$ to 0, attitude $\mathbf{q}$ to the unit quaternion, and covariance matrix $\mathbf{P}$ to 0 except for the following sub-blocks:
-$\mathbf{P}_{b_{a}\,b_{a}}=\mathbf{I}\eta^{2}_{b_{a}};\;\mathbf{P}_{b_{\omega}\,b_{\omega}}=\mathbf{I}\eta^{2}_{b_{\omega}};\;\mathbf{P}_{\delta\theta_{xy}\,\delta\theta_{xy}}=\mathbf{I}\eta^{2}_{\theta},$ (8)
-where $\eta_{b_{a}}$, $\eta_{b_{\omega}}$ and $\eta_{\theta}$ are the initial uncertainties (standard deviations) of the accelerometer bias, gyroscope bias and roll/pitch attitude error.
-We consider the initial yaw attitude uncertainty to be zero because we are not incorporating a compass/magnetometer.
+First, we initialize all components of the state vector $\mathbf{x}$ to 0, and the rotational variables to the unit quaternion.
+The initial radar-to-IMU frame ($\mathbf{t}_{r}^{b}$, $\mathbf{q}_{r}^{b}$) can be supplied externally.
+The robot is assumed to remain still during the first few seconds, which allows for averaging initial accelerometer/gyroscope readings and extracting an estimate of the body-frame gravity vector $\tilde{\mathbf{g}}_{b}$ (after flipping sign) and the gyroscope bias $\mathbf{b}_{\omega}$ respectively.
+An initial value for the roll/pitch components of $\mathbf{q}_{b}^{w}$ can then be derived from $\tilde{\mathbf{g}}_{b}$, while the yaw component is set to 0 since we are not incorporating a compass/magnetometer.
+In addition, we derive the initial accelerometer bias as $\mathbf{b}_{a}=\mathbf{C}_{w}^{b}\left(\mathbf{C}_{b}^{w}\tilde{\mathbf{g}}_{b}-\mathbf{g}\right)$, thus taking into account the difference in magnitude between the expected and observed gravity vectors.
+We initialize the covariance matrix mostly as 0, except for the subblocks related to $\mathbf{t_{r}^{b}}$, $\mathbf{b}_{a}$, $\mathbf{b}_{\omega}$, $\delta\boldsymbol{\theta}_{b}^{w}$ and $\delta\boldsymbol{\theta}_{r}^{b}$, whose diagonals we fill in according to supplied initial values.
 V-B EKF propagation
@@
 These three matrices are sparsely populated: $\mathbf{F}=\mathbf{I}$ and $\mathbf{Q}=\mathbf{N}=\textbf{0}$, except for the following subblocks:
-$\mathbf{F}_{p\,v}=\mathbf{I}t;\;\mathbf{F}_{p\,b_{a}}=-\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{F}_{p\,\delta\theta}=-\frac{1}{2}[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t^{2};$ (13)
-$\mathbf{F}_{v\,b_{a}}=\mathbf{F}_{\delta\theta\,b_{\omega}}=-\mathbf{C}^{w}_{b}t;\;\mathbf{F}_{v\,\delta\theta}=-[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t;$ (14)
+$\mathbf{F}_{p\,v}=\mathbf{I}t;\;\mathbf{F}_{p\,b_{a}}=-\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{F}_{p\,\delta\theta_{b}^{w}}=-\frac{1}{2}[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t^{2};$ (8)
+$\mathbf{F}_{v\,b_{a}}=\mathbf{F}_{\delta\theta_{b}^{w}\,b_{\omega}}=-\mathbf{C}^{w}_{b}t;\;\mathbf{F}_{v\,\delta\theta_{b}^{w}}=-[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t;$ (9)
 $\mathbf{Q}_{v\,v}=\mathbf{I}\sigma^{2}_{v};\;\mathbf{Q}_{\theta\,\theta}=\mathbf{I}\sigma^{2}_{\theta};\mathbf{Q}_{a\,a}=\mathbf{I}\frac{\sigma^{2}_{a}}{t};\;$ (10)
 $\mathbf{Q}_{\omega\,\omega}=\mathbf{I}\frac{\sigma^{2}_{\omega}}{t};\mathbf{Q}_{b_{a}\,b_{a}}=\mathbf{I}\sigma^{2}_{b_{a}}t;\;\mathbf{Q}_{b_{\omega}\,b_{\omega}}=\mathbf{I}\sigma^{2}_{b_{\omega}}t;$ (11)
-$\mathbf{N}_{p\,a}=\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{N}_{v\,a}=\mathbf{N}_{\delta\theta\,\omega}=\mathbf{C}^{w}_{b}t;$ (17)
-$\mathbf{N}_{v\,v}=\mathbf{N}_{\delta\theta\,\theta}=\mathbf{N}_{b_{a}\,b_{a}}=\mathbf{N}_{b_{\omega}\,b_{\omega}}=\mathbf{I},$ (18)
+$\mathbf{N}_{p\,a}=\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{N}_{v\,a}=\mathbf{N}_{\delta\theta_{b}^{w}\,\omega}=\mathbf{C}^{w}_{b}t;$ (12)
+$\mathbf{N}_{v\,v}=\mathbf{N}_{\delta\theta_{b}^{w}\,\theta}=\mathbf{N}_{b_{a}\,b_{a}}=\mathbf{N}_{b_{\omega}\,b_{\omega}}=\mathbf{I},$ (13)
 where $\sigma_{v}$, $\sigma_{\theta}$, $\sigma_{a}$, $\sigma_{\omega}$, $\sigma_{b_{a}}$ and $\sigma_{b_{\omega}}$ are the standard deviations of the respective components of $\mathbf{w}$.
@@
 $\mathbf{x}\leftarrow\mathbf{x}+\mathbf{K}\mathbf{r};\;\mathbf{P}\leftarrow\mathbf{L}\mathbf{P}\mathbf{L}^{T}+\mathbf{K}\mathbf{R}\mathbf{K}^{T}.$ (15)
-As previously mentioned, the attitude part of the robot’s pose is expressed as error state in the EKF, and after an EKF update the nominal value of $\delta\boldsymbol{\theta}$ becomes non-zero, making it necessary to incorporate the estimated error back into the state as follows:
+As previously mentioned, rotations are expressed as error state in the EKF, and after an EKF update the nominal value of each $\delta\boldsymbol{\theta}$ becomes non-zero, making it necessary to incorporate the estimated error back into the state as follows:
 $\mathbf{q}\leftarrow\delta\mathbf{q}\otimes\mathbf{q};\;\delta\mathbf{q}=\exp\left(\frac{1}{2}\delta\boldsymbol{\theta}\right),$ (16)
 $\delta\boldsymbol{\theta}\leftarrow 0;\;\mathbf{P}\leftarrow\mathbf{G}\mathbf{P}\mathbf{G}^{T};\;\mathbf{G}_{\delta\theta\,\delta\theta}=\mathbf{R}\{\delta\mathbf{q}\},$ (17)
-where $\mathbf{G}$ is the error reset matrix, and it is a sparse matrix equal to $\mathbf{I}$ except for the given subblock.
+where $\mathbf{G}$ is the error reset matrix, and it is a sparse matrix equal to $\mathbf{I}$ except for every $\delta\theta\,\delta\theta$ subblock.
+Finally, in order to prevent especially noisy observations from causing instability in the filter, we perform Mahalanobis outlier rejection, using a threshold calculated from a given percentile of the $\chi^{2}$ distribution, like [4].
 V-D Radar egovelocity observation
 Following previous works, we directly observe the robot’s velocity by leveraging the Doppler effect in measurements returned by the radar.
-The returned point cloud is first transformed by $\mathbf{C}_{r}^{b}$ to match the IMU orientation (we consider the separation between IMU and radar to be negligible).
-Afterwards, we use the RANSAC-LSQ algorithm [3] to both filter out outlier (dynamic) points, and estimate the current body-frame velocity of the robot, including a covariance matrix.
+We use the RANSAC-LSQ algorithm [4] to both filter out outlier (dynamic) points, and estimate the current radar-frame velocity of the robot, including a covariance matrix.
 The filtered point cloud is also fed to the Gaussian modeling and scan matching algorithm in the following subsection.
 The observation model used to perform the EKF update is as follows:
-$\mathbf{r}=\mathbf{y}-\mathbf{C}^{b}_{w}\mathbf{v};\;\mathbf{H}_{:\,v}=\mathbf{C}^{b}_{w};\;\mathbf{H}_{:\,\delta\theta}=\mathbf{C}^{b}_{w}[\mathbf{v}]_{\times}.$ (23)
+$\mathbf{r}=\mathbf{y}-{\mathbf{C}_{b}^{r}\left([\boldsymbol{\omega}]_{\times}\mathbf{t}_{r}^{b}+\mathbf{C}^{b}_{w}\mathbf{v}\right)};$ (18)
+$\mathbf{H}_{:\,v}={\mathbf{C}_{b}^{r}}\mathbf{C}_{w}^{b};\;{\mathbf{H}_{:\,t_{r}^{b}}=\mathbf{C}_{b}^{r}[\boldsymbol{\omega}]_{\times}};\;{\mathbf{H}_{:\,b_{\omega}}=\mathbf{C}_{b}^{r}[\mathbf{t}_{r}^{b}]_{\times}};$ (19)
+$\mathbf{H}_{:\,\delta\theta_{b}^{w}}={\mathbf{C}_{b}^{r}}\mathbf{C}_{w}^{b}[\mathbf{v}]_{\times};\;{\mathbf{H}_{:\,\delta\theta_{r}^{b}}=\mathbf{C}_{b}^{r}[[\boldsymbol{\omega}]_{\times}\mathbf{t}_{r}^{b}+\mathbf{C}^{b}_{w}\mathbf{v}]_{\times}}.$ (20)
 V-E Scan matching observation
@@
 We are currently using a fixed observation covariance matrix $\mathbf{R}$, although we believe it should be possible to leverage the optimized particle swarm returned by the scan matching algorithm to estimate this matrix.
-Furthermore, we constrain the observation model described above by removing degrees of freedom that are deemed noisy in the scan matching due to the properties of the sensor.
-For example, radars have lower resolution in the Z axis, which degrades the usefulness of Z/roll/pitch information returned by scan matching algorithms.
-In order to ignore the unreliable degrees of freedom, we transform the residual vector $\mathbf{H}_{c}\mathbf{r}$, observation matrix $\mathbf{H}_{c}\mathbf{H}$ and covariance matrix $\mathbf{H}_{c}\mathbf{R}\mathbf{H}_{c}^{T}$ by a constraint matrix defined as follows:
-$\mathbf{H}_{c}=\begin{bmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&1\end{bmatrix}.$ (24)
-Table I: Quantitative Analysis on NTU4DRadLM Dataset using EVO. Best results in bold, second best results underlined
+Table I: Results on NTU4DRadLM Dataset (Oculii Eagle). Best results in bold, second best results underlined.
+Methods marked with * are self-reported metrics. EV: Egovelocity, SM: Scan Matching
 cp nyl loop2 loop3
-$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
-(%) (º/m) (%) (º/m) (%) (º/m) (%) (º/m)
-NTU4DRadLM baseline: gicp [25] 4.87 0.0648 5.65 0.0257 6.73 0.0094 6.03 0.0097
-NTU4DRadLM baseline: Fast-LIO [25] 3.97 0.0773 4.47 0.0273 8.38 0.0096 6.95 0.0105
-Our method (GICP scan matching) 1.69 0.0323 4.50 0.0198 3.92 0.0044 3.81 0.0054
-Our method (single hypothesis) 1.47 0.0284 3.17 0.0150 2.95 0.0043 4.80 0.0065
-Our method (full version) 1.33 0.0260 2.92 0.0134 3.49 0.0047 4.55 0.0061
-Table II: Quantitative Analysis on NTU4DRadLM Dataset using rpg_trajectory_evaluation. Methods marked with * are self-reported metrics
-cp nyl loop2 loop3
-$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
-(%) (º/m) (%) (º/m) (%) (º/m) (%) (º/m)
+Usage $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
+IMU EV SM (%) (º/m) (%) (º/m) (%) (º/m) (%) (º/m)
 NTU4DRadLM baseline: gicp* [29] 4.13 0.0552 4.62 0.0184 4.84 0.0060 3.22 0.0060
@@
 RIV-SLAM* [24] 2.58 0.0342 2.69 0.0456
-NTU4DRadLM baseline: gicp [25] 4.921 0.068 5.775 0.027 6.743 0.010 6.242 0.010
-NTU4DRadLM baseline: Fast-LIO [25] 4.178 0.083 4.759 0.030 8.221 0.010 7.120 0.011
-Our method (GICP scan matching) 1.746 0.035 4.142 0.020 3.902 0.005 3.781 0.006
-Our method (single hypothesis) 1.501 0.030 3.133 0.015 2.851 0.004 4.529 0.007
-Our method (full version) 1.383 0.028 2.934 0.014 3.413 0.005 4.259 0.006
-Figure 3: Generated trajectories for the 4 tested NTU4DRadLM sequences.
+EKF-RIO [4] 14.76 0.3955 35.89 0.2464 44.31 0.0349 50.23 0.0910
+(Ablation) Our EKF with NDT [2] 2.42 0.0345 7.80 0.0408 5.61 0.0065 7.84 0.0100
+(Ablation) Our EKF with GICP [21] 2.93 0.0428 2.97 0.0111 2.99 0.0038 3.50 0.0042
+(Ablation) Our EKF with VGICP [12] 2.43 0.0362 2.56 0.0136 2.73 0.0036 3.84 0.0047
+(Ablation) Gaussian single hypothesis 2.15 0.0318 2.46 0.0116 2.44 0.0030 3.97 0.0043
+Gaussian multi-hypothesis 1.64 0.0310 2.41 0.0117 2.61 0.0036 3.71 0.0040
+Table II: Results on Snail-Radar Dataset (Oculii Eagle). Best results in bold, second best results underlined.
+st_20231213_1 iaf_20231213_2 iaf_20231213_3 if_20231213_4 if_20231213_5
+$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
+(%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m)
+(Ablation) Our EKF with NDT [2] 4.13 0.5286 2.04 0.0519 3.95 0.0653 2.86 0.0786 3.15 0.0771
+(Ablation) Our EKF with GICP [21] 3.77 0.5118 2.93 0.0629 4.05 0.0629 2.77 0.1136 4.58 0.1572
+(Ablation) Our EKF with VGICP [12] 3.68 0.5141 3.06 0.0610 3.17 0.0551 2.53 0.1167 4.44 0.1489
+(Ablation) Gaussian single hypothesis 3.36 0.5204 2.85 0.0689 2.84 0.0537 2.36 0.0927 4.00 0.1457
+Gaussian multi-hypothesis 3.33 0.5147 2.45 0.0543 2.89 0.0498 2.55 0.1093 4.32 0.1569
+Table III: Results on Snail-Radar Dataset (ARS548). Best results in bold, second best results underlined.
+st_20231213_1 iaf_20231213_2 iaf_20231213_3 if_20231213_4 if_20231213_5
+$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
+(%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m)
+(Ablation) Our EKF with GICP [21] 1.69 0.3049 3.36 0.0507 3.29 0.0487 2.42 0.0761 2.26 0.0796
+(Ablation) Our EKF with VGICP [12] 1.95 0.2798 2.57 0.0494 3.10 0.0491 2.81 0.0777 2.12 0.0836
+(Ablation) Gaussian single hypothesis 1.34 0.3727 2.50 0.0487 2.77 0.0475 2.47 0.0826 2.37 0.0740
+Gaussian multi-hypothesis 1.60 0.3573 4.02 0.0756 6.80 0.1000 2.93 0.0894 2.25 0.0746
+Furthermore, we constrain the observation model described above by removing degrees of freedom that are deemed noisy in the scan matching due to the properties of the sensor.
+For example, radars have lower resolution in the Z axis, which degrades the usefulness of Z/roll/pitch information returned by scan matching algorithms.
+In order to ignore the unreliable degrees of freedom, we transform the residual vector, observation matrix and covariance matrix by removing the rows and/or columns corresponding to Z/roll/pitch prior to applying the EKF update.
 VI Experimental results
-We compare our odometry with previous published 4D radar-inertial odometry methods on the NTU4DRadLM dataset [25].
+In order to evaluate the performance of our Gaussian modeling and scan matching method, we use existing public 4D radar datasets.
+In particular, we selected the NTU4DRadLM [29] and Snail-Radar [8] datasets, because together they cover multiple radar types (Oculii Eagle and ARS548), multiple robot platform types (handcart and car), and adverse conditions.
+We evaluate relative translation/rotation errors using the modern and flexible evo evaluation package [7], while preserving existing practice by using a custom evaluation script that follows the behavior of older packages such as rpg_trajectory_evaluation as closely as possible.
+In the case of Snail-Radar sequences, we report rotation errors in mrad/m instead of º/m because the errors are much smaller.
+Besides our full method, we evaluate four ablations within our EKF framework: one without multi-hypothesis registration, and three replacing our Gaussian-based algorithms with NDT, GICP and VGICP.
+Since these are all single-hypothesis methods, we use the current nominal state of the EKF directly as the initial estimate (instead of sampling a random particle swarm).
+The NDT version uses the implementation provided by the well known PCL (Point Cloud Library), while the GICP and VGICP versions use small_gicp [11] as the underlying implementation.
+In the case of Oculii Eagle scans, we set a target of 20 points per Gaussian; while for ARS548 scans, due to the much lower number of returned points, we set a target of 8 points per Gaussian.
+In both cases we sample a swarm of 8 particles during scan matching.
+The source code of our odometry solution can be accessed online on GitHub222 https://github.com/robotics-upo/gaussian-rio-cpp .
+We include the script used to generate the odometry trajectories, and the script used to evaluate trajectories with evo.
+The code also contains all details regarding the system parameters we used, such as EKF covariance initialization, IMU/noise parameters, etc.
+VI-A Results on NTU4DRadLM
+We compare our odometry with previous published results on the NTU4DRadLM dataset [29], which includes scans captured by an Oculii Eagle radar, and a VectorNav VN-100 IMU.
 These include the GICP and Fast-LIO baselines tested in the original NTU4DRadLM paper, the APDGICP algorithm tested in the follow-up 4DRadarSLAM paper [30], as well as other works such as EFEAR-4D [26] or RIV-SLAM [24].
-We compare versions of these methods without loop closure if possible, in order to focus on evaluating pure odometry.
-Out of the six sequences present in NTU4DRadLM, we perform evaluation using four: cp and nyl (low speed regime, using a handcart platform), and loop2 and loop3 (high speed regime, using a car platform).
-The garden and loop1 sequences cannot be reliably used for evaluation because they contain large gaps in the IMU and radar data, respectively.
-We evaluate relative translation/rotation errors using the modern and flexible evo evaluation package [6], as well as the older rpg_trajectory_evaluation package [27] used by previous works, using default settings.
-In the case of evo we use a custom evaluation script written specifically to follow the default behavior of rpg_trajectory_evaluation as closely as possible.
-We run both evaluation packages on the baseline GICP and Fast-LIO trajectories published along with the NTU4DRadLM dataset, as well as the trajectories generated by our odometry solution.
-In particular, we report results for the full version of our method, and also two ablated versions: one that replaces the Gaussian modeling and scan matching with GICP, and another that uses single-hypothesis scan matching.
-The GICP version uses small_gicp [9] as the underlying implementation, and in both ablations we use the current nominal state of the EKF directly as the initial estimate (instead of sampling a random particle swarm).
-We also include the self-reported metrics for all methods that appear in their respective papers as-is, since the trajectory files needed for metric calculation are not available.
-The source code of our odometry solution can be accessed online on GitHub [^2] https://github.com/robotics-upo/gaussian-rio.
-We include the script used to generate the odometry trajectories, and the script used to evaluate trajectories with evo.
-The code also contains all details regarding the system parameters we used, such as EKF covariance initialization, IMU/noise parameters, number of Gaussians, etc.
-VI-A Quantitative Analysis
-VI-A1 Evaluation with evo
-In Table I, we can observe major improvements in all 4 sequences provided by our method compared to the baselines originally tested by the authors of NTU4DRadLM.
-Moreover, the comparison between the Gaussian versions of our method and our GICP ablation shows improvements in relative translation error ($t_{rel}$) in 3 out of 4 sequences, and improvements in relative rotation error ($r_{rel}$) in 3 out of 4 sequences.
-In particular, the best improvements can be seen in the nyl sequence.
-The version with scan matching based on multiple hypotheses overperforms the version with a single hypothesis in 3 out of 4 sequences.
-We theorize that the sequences where our GICP ablation overperforms Gaussian versions of our method contain more challenging/sparser geometry due to being higher speed car driving sequences in open environments.
-Nonetheless, we believe it should be possible to further adjust the hyperparameters used (such as number of Gaussians used for modeling, number of hypotheses used for scan matching, keyframe thresholds, EKF covariances used, etc.) in order to optimize the algorithm for those sequences.
-VI-A2 Evaluation with rpg_trajectory_evaluation
-In Table II, we note discrepancies in the NTU4DRadLM baseline metrics between the reported values in the paper and the values obtained by running the evaluation software on the corresponding published trajectories, especially in the higher speed sequences.
-We also note very similar results for all methods generated by both evaluation packages, validating the use of both packages for evaluating 4D radar-inertial odometry methods.
-The errors reported by EFEAR-4D for the higher speed sequences are much higher than the values reported by other methods.
-We believe this to be caused by some difference in the evaluation methodology, as the same paper reports similar errors for baseline methods such as [26].
-The most challenging sequence is loop3, and interestingly the best results are the ones reported by the NTU4DRadLM authors for the simplest baseline method.
-RIV-SLAM is the best performing method in loop2, however it must be said that to the best of our knowledge this method includes loop closure.
-The second best performing method in that sequence is the single-hypothesis version of our method, in line with the results in Table I.
-VI-B Qualitative Analysis
-Figure 3 shows the plotted trajectories of the ground truth, NTU4DRadLM GICP baseline and the full version of our method; across the 4 sequences.
-We can observe that in the low speed sequences (cp, nyl) our method accumulates substantially less drift in the XY plane, and manages to closely match the ground truth.
-High speed sequences (loop2, loop3) prove to be more challenging for both methods; in particular, our method drifts the most in loop3 with respect to the ground truth.
-On the other hand, our method performs much better than the baseline in the XZ plane across all 4 sequences.
-In particular, in loop2 and loop3 our method avoids large drifts in the Z axis, compared to the drifts in the baseline of over 70 meters.
-We think the higher accuracy in the XZ plane for our methods is responsible for improved quantitative results in these sequences, despite the lower accuracy in the XY plane.
+We include these self-reported metrics as-is, as the trajectory files needed for metric calculation are usually not made public.
+Even though there are no published results for EKF-RIO [4] on this dataset, it is one of the few methods with available source code, and so we are able to produce and include evaluation results for it.
+We report metrics without loop closure whenever possible, in order to focus on evaluating pure odometry.
+We evaluate the cp, nyl (low speed handcart), loop2 and loop3 (high speed car) sequences.
+The garden and loop1 sequences are excluded because they contain large gaps in the IMU and radar data, respectively.
+Table I shows the comparison between all methods.
+We can observe how our Gaussian methods provide the best or second best results for all sequences except for loop3, which seems to favor GICP-based methods (both baseline and our implementation).
+Moreover, in 3 out of the 4 sequences the multi-hypothesis scan matching approach provides improvements in translation error, especially in cp.
+VI-B Results on Snail-Radar
+In addition to NTU4DRadLM, we perform additional testing on the Snail-Radar dataset [8], which uses both Oculii Eagle and ARS548 radars together with an Xsens MTi-3-DK IMU, and includes sequences recorded in harsher environment conditions.
+In particular, we selected the five sequences recorded on 2023/12/13 because of the nighttime and rain conditions of that day under a high speed car platform.
+These sequences are all fairly long, with the longest (iaf_20231213_2) being 15 min and most being around 7 min in length.
+Table II shows results using the Oculii Eagle radar.
+In general, we can observe the best results to be spread across the different methods.
+Nonetheless, our Gaussian algorithms obtain the best or second best results in 9 out of 10 metrics, and the best translation error in 3 out of 5 sequences (the other two favoring NDT).
+The multi-hypothesis approach improves or matches results compared to the single hypothesis version in 3 out of 5 sequences.
+It is worth mentioning that NDT, unlike in NTU4DRadLM, seems to behave much better in Snail-Radar.
+Table III shows results using the ARS548 radar.
+This radar produces much sparser point clouds than Oculii Eagle (around 10x fewer points).
+We believe this to be caused by the radar lacking the scan enhancement algorithm used by Oculii Eagle.
+However interestingly, we can observe better metrics in general compared to Oculii Eagle.
+The sparser scans trigger a known issue in PCL’s implementation of NDT that cause the program to abnormally terminate, so unfortunately NDT is excluded from the comparison.
+The results once again show different methods leading in different sequences, however it is worth pointing out that 6 out of the 10 best results are now obtained by our Gaussian algorithms, as well as 3 out of the 10 second best results.
+Another thing to mention is that the multi-hypothesis scan matching seems to perform worse than the single-hypothesis version.
+We hypothesize this to be related to the increased sparsity of the scans, and believe it could be solved by improving the sampling distribution of the particles.
 VI-C Summary of results
 We consider that the results described above validate our methodology.
-Our methods produce better relative translation and rotation metrics in 3 out of 4 tested sequences.
-Moreover, they produce generally better results regarding Z axis drift.
-Our ablation experiments also validate the Gaussian-based modeling and scan matching algorithm, as well as the multi-hypothesis approach; which contribute to improve both estimation accuracy and robustness.
+Our Gaussian modeling and scan matching algorithm produces comparable relative translation and rotation metrics to existing registration algorithms, as shown by the experiments.
+Moreover, in several sequences across several datasets and radar types our algorithm produces the best overall results.
+Our EKF-based radar-inertial framework is also validated by the experiments, which show how on a publicly evaluated dataset we are able to improve on the state of the art even when not using our Gaussian algorithms.
+On the other hand, the multi-hypothesis approach seems to need further adjustments, although it is already capable of producing better results for some sequences.
+As an additional observation, it is worth mentioning that we believe the radar-to-IMU calibration file provided by NTU4DRadLM to be inaccurate, as it does not seem to match the placement of the different sensors in the photo they provide of their platform.
+The effects of this inaccurate calibration can be especially seen in the rotational errors, which are much higher for all methods compared to those achieved under Snail-Radar sequences.
 VII Conclusion
-In this paper we introduce a multi-hypothesis Gaussian-based radar-inertial odometry pipeline.
+In this paper we introduce a Gaussian-based radar-inertial odometry pipeline.
 Our methods summarize noisy 4D radar scans using freely positioned 3D Gaussians (like 3DGS [10] and unlike NDT [2]), which then support multi-hypothesis scan matching for increased robustness.
-We show that the fusion of these methods into an EKF produces better quantitative and qualitative results than existing baselines using traditional scan matching algorithms.
-Future works include improving the Gaussian modeling and scan-matching algorithms (in particular, exploring optimization algorithms more direct than gradient descent), modeling kinematic restrictions relevant to the robot platform used in order to reduce noise during egovelocity estimation, further refining the pose estimates by means of Pose Graph Optimization (PGO), incorporating a loop closure module, and creating a global Gaussian map of the environment for the purposes of full SLAM.
+Through experiments in multiple datasets with multiple radar types, we show that these methods produce results comparable to existing registration algorithms, and even manage to outperform them in several sequences.
+Future works include improving the Gaussian models with an optimization algorithm more efficient than gradient descent, better initialization and dynamic management of Gaussians (as in 3DGS), as well as incorporating radar-specific information such as intensity values; improving the robustness of the multi-hypothesis particle system with kinematic priors, dynamically adjusting the hyperparameters of the system, improving the quality of egovelocity estimation by modeling kinematic restrictions relevant to the robot platform used, further refining the pose estimates by means of Pose Graph Optimization (PGO), incorporating a loop closure module, and creating a global Gaussian map of the environment for the purposes of full SLAM.
 Acknowledgment

### v3

@@
 Abstract
-4D millimeter-wave (mmWave) radars are sensors that provide robustness against adverse weather conditions (rain, snow, fog, etc.), and as such they are increasingly being used for odometry and SLAM applications.
-However, the noisy and sparse nature of the returned scan data proves to be a challenging obstacle for existing point cloud matching based solutions, especially those originally intended for more accurate sensors such as LiDAR.
-Inspired by visual odometry research around 3D Gaussian Splatting [10], in this paper we propose using freely positioned 3D Gaussians to create a summarized representation of a radar point cloud tolerant to sensor noise, and subsequently leverage its inherent probability distribution function for registration (similar to NDT [2]).
-Moreover, we propose simultaneously optimizing multiple scan matching hypotheses in order to further increase the robustness of the system against local optima of the function.
-Finally, we fuse our Gaussian modeling and scan matching algorithms into an EKF radar-inertial odometry system designed after current best practices.
-Experiments using publicly available 4D radar datasets show that our Gaussian-based odometry is comparable to existing registration algorithms, outperforming them in several sequences.
+4D millimeter-wave (mmWave) radars are sensors that provide robustness against adverse weather conditions (rain, snow, fog, etc.), and as such they are increasingly used for odometry and SLAM (Simultaneous Location and Mapping).
+However, the noisy and sparse nature of the returned scan data proves to be a challenging obstacle for existing registration algorithms, especially those originally intended for more accurate sensors such as LiDAR.
+Following the success of 3D Gaussian Splatting for vision, in this paper we propose a summarized representation for radar scenes based on global simultaneous optimization of 3D Gaussians as opposed to voxel-based approaches, and leveraging its inherent probability distribution function for registration.
+Moreover, we propose tackling the problem of radar noise by optimizing multiple scan matching hypotheses in order to further increase the robustness of the system against local optima of the function.
+Finally, following existing practice we implement an Extended Kalman Filter-based Radar-Inertial Odometry pipeline in order to evaluate the effectiveness of our system.
+Experiments using publicly available 4D radar datasets show that our Gaussian approach is comparable to existing registration algorithms, outperforming them in several sequences.
 Our code and results can be publicly accessed at: https://github.com/robotics-upo/gaussian-rio-cpp
@@
 As such, there is considerable interest in incorporating these sensors in odometry and SLAM applications.
-Existing LiDAR odometry solutions have been applied, adapted and sometimes even further specialized to 4D radar, with varying degrees of success [28, 29, 30, 26, 24].
-However, the point clouds generated by 4D radars are sparser, noisier, and have a narrower field of view compared to those generated by LiDARs; and as a result these algorithms often struggle with the geometry of certain scenes.
-Therefore, it is desirable to explore scene representations that directly model coarse-grained surfaces and regions of radar point clouds instead of focusing on individual points themselves.
-Another technique that has gained popularity recently in the field of robotics is 3D Gaussian Splatting (3DGS) [10].
-This algorithm creates an explicit representation of a visual scene, and has been applied to SLAM [16, 9, 1] as an alternative to existing methods using implicit representations such as NeRF [17].
-However, 3DGS heavily relies on the nature of visual information for training and rendering, and thus adapting it for use with LiDAR or radar instead of visual cameras is challenging; as opposed to NeRF, which can be straightforwardly adapted to model a SDF instead [18].
+A key component of such system is the modeling and scan matching algorithm.
+Point clouds are used to capture the geometry of the scene at a given time, and subsequent scans are compared (matched) to the reference model in order to extract localization information.
+Popular established algorithms that perform this task include the Normal Distributions Transform (NDT) [1] or Generalized Iterative Closest Point (GICP) [2].
+However, despite working well for LiDAR [3, 4, 5], radar point clouds are considerably more challenging to register because they are sparser, noisier, and have a narrower field of view; and, as a result, these algorithms often struggle with the geometry of certain scenes.
 Figure 1: Gaussian model of an example point cloud, containing 150 Gaussians. The surface of each spheroid represents the $2\sigma$ front of each Gaussian (i.e. Mahalanobis distance equal to two).
-In this paper we draw inspiration from both lines of research to propose the use of Gaussian models for representing a point cloud and performing registration with respect to it, as an alternative to existing registration methods such as the Normal Distributions Transform (NDT) [2] or Generalized Iterative Closest Point (GICP) [21].
-Gaussians are freely positioned in space instead of following a regular grid, as in 3DGS and unlike NDT.
-Moreover, our scan matching (registration) step considers multiple simultaneous hypotheses, which can in some situations increase the performance of the system, as we show in our experiments.
-In addition, we outline an Extended Kalman Filter (EKF) based approach to radar-inertial odometry that consolidates existing practices by integrating both radar egovelocity [4] and keyframing/scan matching components, which we then use to evaluate our Gaussian-based approach in comparison to other registration algorithms on multiple datasets and radar types.
-Figure 2: Overview of the proposed Gaussian radar-inertial odometry pipeline
-Figure 2 shows the overall pipeline of our odometry system.
-As in other solutions, we use an Extended Kalman Filter (EKF) to perform inertial strapdown and process sensory input, including radar egovelocity.
-Keyframing is used to select individual 4D radar scans, which then undergo Gaussian modeling.
-The Gaussian multi-hypothesis scan matching algorithm registers every incoming 4D radar scan against the Gaussian model in order to produce a relative pose observation with respect to the last keyframe, which is then used to further update the EKF.
-This paper is organized as follows: Section II reviews existing related literature, Sections III, IV and V describe the three main components of our method respectively: Gaussian modeling, Gaussian scan matching, and the Extended Kalman Filter tying everything together; Section VI shows experimental results, and Section VII closes the paper with some conclusions.
+Figure 2: Overview of the Radar-Inertial Odometry pipeline including our proposed Gaussian modeling and scan matching processes
+We introduce a new method for modeling and registering 4D radar point clouds using 3D Gaussians.
+Following the success of 3D Gaussian Splatting (3DGS) for vision [6], we propose the use of a collection of 3D Gaussians for modeling the geometry of a scene (see Fig.
+1).
+As in [6] (and in contrast to NDT [1]), we simultaneously optimize all parameters of the model, including the position, scaling, rotation and correspondences between points and Gaussians.
+This approach is entirely featureless, and does not require any priors other than an initial set of centroids.
+Moreover, we propose a registration step that considers multiple simultaneous hypotheses in order to add further robustness against radar noise.
+Finally, we integrate our algorithms into a Radar-Inertial Odometry (RIO) pipeline using an Extended Kalman Filter (EKF) (see Fig.
+2), which we then use to evaluate our Gaussian methods against several established registration algorithms on multiple datasets and radar types.
+This paper is organized as follows: Section II reviews existing related literature, Sections III and IV describe the proposed Gaussian modeling and scan matching components respectively, Section V describes the EKF-based RIO pipeline used in the experiments; Section VI shows experimental results, and Section VII concludes the paper proposing future work.
 II Related work
-II-A Odometry and SLAM
-There exist a large number of works focused on odometry, sometimes even attaining a full SLAM pipeline (including loop closure).
-Particularly we are interested in approaches that use non-visual sensors, such as [27, 25, 19, 4, 5, 28, 29, 30, 26, 24, 31].
-These works use a 3D LiDAR or 4D radar sensor, sometimes together with an IMU, to provide an odometry solution.
-The main technical advantage of these sensors over visual sensors (cameras) is their increased field/depth of view and robustness to adverse environmental conditions: in particular, both are immune to low light conditions, and 4D radar in addition offers robustness to fog, dust and rain.
-Moreover, 4D radar is uniquely capable of measuring the velocity of objects by leveraging the Doppler effect.
-Researchers have taken advantage of this to directly measure the robot’s own velocity (“egovelocity”) [4, 5], something which has been shown to significantly improve the performance of odometry systems.
-Registration of incoming point clouds with respect to previous ones is a key technique used in some of the cited odometry systems.
+Registration of incoming point clouds with respect to previous ones is a key technique used in 3D LiDAR and 4D radar odometry and SLAM systems [3, 4, 5, 7, 8, 9, 10, 11, 12].
 In particular, the most widely used algorithm for aligning point clouds is Generalized Iterative Closest Point (GICP) [2].
-This algorithm calculates distribution probability at each point based on their neighborhood, and optimizes a distribution-to-distribution scan matching function.
+This algorithm calculates a distribution probability at each point based on its neighborhood, and optimizes a distribution-to-distribution scan matching function.
 In the case of radar scans, this is a more difficult task compared to LiDAR because of the lower number and density of points, as well as higher measurement noise/uncertainty in each point’s position.
@@
 However, this is still in essence an approach focused on individual points, without attempting to extract larger scale geometric features that may be more stable and robust for registration purposes.
-The works cited in this section use different sources of information, however many do not leverage all mentioned sources at the same time. Table I includes a list of surveyed methods.
-II-B Gaussian-based modeling
-The Normal Distributions Transform (NDT) algorithm [2] slices and subdivides a point cloud using a regular grid and computes normal distributions for each cell that contains points.
+The Normal Distributions Transform (NDT) algorithm [1] subdivides a point cloud using a voxel grid and computes normal distributions for each cell that contains points.
 Scan matching is performed by maximizing the resulting probability distribution function with respect to an incoming scan.
-An important limitation of this algorithm is the fact that the grid structure itself imposes restrictions on the geometry, and the resulting model may not be fine grained enough to accurately model the geometry of the environment.
-Another problem lies in the fact that discontinuities in the probability function are produced when partially registered points cross cell boundaries during optimization.
-Moreover, the original authors only consider 2D scan matching (requiring 3D point clouds to be projected into the XY plane), although subsequent works [15] extend the algorithm to 3D.
-NDT has been used with radar scans too [13, 14], albeit without inertial information.
-Another similar algorithm is the Voxelized GICP (VGICP) [12], which also creates an NDT-like voxel-based representation to serve as the reference model, and combines it with the distribution-to-distribution based approach of GICP.
+An important limitation of this algorithm is that the grid structure itself imposes restrictions on the geometry, and the resulting model may not be fine grained enough to accurately model the geometry of the environment.
+Another problem lies in that discontinuities in the probability function are produced when partially registered points cross cell boundaries during optimization.
+NDT has been used with radar scans too [13, 14].
+Voxelized GICP (VGICP) [15] is a derivative that combines NDT’s voxel-based representation with the distribution-to-distribution based approach of GICP for registration.
+G3Reg [16] is another derivative of NDT, which performs classification on each voxel (plane/cluster/point) in order to merge plane cells together, and extract semantic information to further aid scan matching.
+In all these algorithms, each individual Gaussian is modeled separately using their associated subcloud extracted from the segmentation.
+However, we are interested in approaches that jointly optimize all Gaussians, looking at the entire point cloud.
 3D Gaussian Splatting (3DGS) [6] is a popular technique for creating a model of a scene from multiple posed images.
 A 3DGS model is a collection of multiple 3D normal distributions (Gaussians) freely distributed in space, each also containing color/radiance information encoded using spherical harmonics.
-Even though this technique was originally developed as an alternative to NeRF [17] for novel view synthesis, researchers have successfully leveraged this representation to develop SLAM solutions [16, 9, 1].
-However, 3DGS and these works are all designed for visual sensors (cameras), and they cannot leverage other sensors such as LiDAR or radar due to the nature of the representation.
-This affects the rendering process in particular: while NeRF’s ray casting framework can be easily adapted to directly predict the output of a SDF [18] (therefore removing the volume rendering component from the system), the differentiable tile rasterizer is a core part of 3DGS that cannot be avoided, neither during training nor inference; thus requiring significant redesign efforts in order to synthesize other types of data such as LiDAR scans [3].
+Even though this technique was originally developed as an alternative to Neural Radiance Fields (NeRF) [17] for novel view synthesis, researchers have successfully leveraged this representation to develop SLAM solutions [18, 19, 20].
+Visual-LiDAR methods that improve the robustness of 3DGS with priors sourced from LiDAR input have also been developed [21, 22, 23].
+However, 3DGS and these works are all designed for visual input, and they cannot be easily reworked to fully replace it with other sensors such as LiDAR or radar due to the nature of the representation.
+This affects the rendering process in particular: while NeRF’s ray casting framework can be easily adapted to directly predict the output of a Signed Distance Field (SDF) [24] (therefore removing the volume rendering component from the system), the differentiable tile rasterizer is a core part of 3DGS that cannot be avoided, neither during training nor inference; thus requiring significant redesign efforts in order to synthesize other types of data such as LiDAR scans [25].
+Following 3DGS as inspiration, we define a purely geometric Gaussian modeling algorithm for radar point clouds based on simultaneous optimization of 3D Gaussians.
+This representation fundamentally differs from NDT in the fact that Gaussians are not bound to a predefined voxel grid, and all parameters of the model are jointly optimized instead of subdividing the clouds and individually generating a Gaussian for each subcloud.
 III Gaussian modeling
 Given a point cloud $\mathcal{P}$ contaning $M$ points $\{\mathbf{p}_{i}|i=0,1,\ldots,M\}$ with $\mathbf{p}_{i}\in\mathbb{R}^{3}$, our goal is creating a summarized representation of its geometry for use with downstream tasks such as mapping or localization.
-Following the line of works such as NDT [2] and 3DGS [10], we decided to model the geometry using normal distributions.
+We model the geometry of a given radar scene using normal distributions, as in [1, 6, 15, 16].
 Figure 1 shows an example of a Gaussian model.
@@
   • The parameters of $\mathcal{G}$ (in other words, all $\theta_{j}$) are updated using the gradient of the loss function.
-We define the following loss function affecting a given $\theta_{j}$, derived from the log PDF of the multivariate normal distribution:
+We define the following loss function affecting a given $\theta_{j}$, derived from the log PDF (Probability Density Function) of the multivariate normal distribution:
 $\mathcal{L}_{j}=\frac{1}{2|G_{j}|}\sum_{\mathbf{p}_{i}\in G_{j}}\hat{\mathbf{p}}_{i}^{T}\hat{\mathbf{p}}_{i}+\sum_{k}^{3}\mathbf{s}_{j_{k}}.$ (1)
@@
 A common technique used in odometry systems is keyframing.
-Certain poses of the robot are selected to become local reference points (keyframes) that are later used for registration, graph optimization or loop closure purposes.
-In our case, we use keyframing to select individual radar point clouds to undergo Gaussian modeling, and serve as reference for Gaussian scan matching of subsequent radar point clouds.
-This has the benefit of reducing the accumulated error that arises from successive applications of the scan matching algorithm, though it also means the criteria used to select keyframe candidates needs to be carefully designed to generate the appropriate keyframe pacing.
+Certain poses of the robot are selected to become local reference points (keyframes) that are typically used for registration, pose graph optimization or (in the case of SLAM) loop closure purposes.
+In our case, we only use keyframing to select individual radar point clouds to undergo Gaussian modeling, and serve as reference for Gaussian scan matching of subsequent radar point clouds.
+The system only retains the pose and Gaussian model of the most recently generated keyframe.
+This approach, compared to trivial successive applications of the scan matching algorithm, has the benefit of reducing the accumulated error, though it also means the criteria used to select keyframe candidates needs to be carefully designed to generate an appropriate keyframe pacing.
 We use a simple set of criteria to decide when a new pose $\xi^{j}$ qualifies to become a new keyframe.
@@
 The criteria are then simply that any of the following must hold:
-  • For the translational component: $||\mathbf{t}||_{2}\geq d_{\max}$, where $d_{\max}$ is the maximum distance threshold.
+  • For the translational component: $\mathbf{t}||_{2}\geq d_{\max}$, where $d_{\max}$ is the maximum distance threshold.
   • For the rotational component: $2\cos^{-1}|q_{w}|\geq\alpha_{\max}$, where $\alpha_{\max}$ is the maximum rotation angle threshold.
@@
 V Inertial odometry with EKF
-We use an Extended Kalman Filter in our odometry solution to tightly integrate all incoming sensor information: inertial data, radar egovelocity observations, and scan matching results.
-Unlike other published solutions [22, 4, 5], we only use Error State filtering for rotational variables, as opposed to the entire state.
+In order to evaluate the suitability and effectiveness of the Gaussian algorithms for Radar-Inertial Odometry, it is necessary to define a pipeline for use in the experiments.
+Following well established practice [28, 29, 12], we use an Extended Kalman Filter in our odometry solution to tightly integrate all incoming sensor information: inertial data, radar egovelocity observations, and scan matching results.
+In order to simplify the formulation, we only use Error State filtering for rotational variables, as opposed to the entire state.
 We employ the right-handed North-West-Up (NWU) convention typical of robotics, and consider three reference frames: world ($w$), body/IMU ($b$) and radar ($r$).
@@
 We follow Kalibr’s IMU noise model [30] [^1] https://github.com/ethz-asl/kalibr/wiki/IMU-Noise-Model, and consider both additive ($\mathbf{w}_{a}$, $\mathbf{w}_{\omega}$) and random walk noise ($\mathbf{w}_{b_{a}}$, $\mathbf{w}_{b_{\omega}}$) in accelerometer/gyroscope readings.
-[…]
+There also exists a vector $\mathbf{g}=-[0\;0\;9.80511]^{T}$ modeling the Earth’s gravity treated as a constant, as we do not consider long enough travel distances for changes in the plane tangent to the Earth’s surface to matter.
 V-A EKF initialization
@@
 Table I: Results on NTU4DRadLM Dataset (Oculii Eagle). Best results in bold, second best results underlined.
-Methods marked with * are self-reported metrics. EV: Egovelocity, SM: Scan Matching
+Methods marked with * are self-reported metrics. EV: Egovelocity, SM: Scan Matching, Abl.: Ablation
 cp nyl loop2 loop3
-Usage $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
-IMU EV SM (%) (º/m) (%) (º/m) (%) (º/m) (%) (º/m)
+Usage Time $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
+IMU EV SM (ms) (%) (º/m) (%) (º/m) (%) (º/m) (%) (º/m)
 NTU4DRadLM Baseline: GICP* [8] 4.13 0.0552 4.62 0.0184 4.84 0.0060 3.22 0.0060
 NTU4DRadLM Baseline: Fast-LIO* [8] 2.94 0.0468 3.80 0.0208 7.16 0.0057 4.55 0.0064
-4DRadarSLAM apdgicp* [30] 3.56 0.0369 3.55 0.0171 6.09 0.0082 4.09 0.0097
+4DRadarSLAM: APDGICP* [9] 3.56 0.0369 3.55 0.0171 6.09 0.0082 4.09 0.0097
 EFEAR-4D* [10] 5.09 0.0125 8.93 0.0166 37.27 0.0152 37.01 0.0175
@@
 EKF-RIO [29] 14.76 0.3955 35.89 0.2464 44.31 0.0349 50.23 0.0910
-(Ablation) Our EKF with NDT [2] 2.42 0.0345 7.80 0.0408 5.61 0.0065 7.84 0.0100
-(Ablation) Our EKF with GICP [21] 2.93 0.0428 2.97 0.0111 2.99 0.0038 3.50 0.0042
-(Ablation) Our EKF with VGICP [12] 2.43 0.0362 2.56 0.0136 2.73 0.0036 3.84 0.0047
-(Ablation) Gaussian single hypothesis 2.15 0.0318 2.46 0.0116 2.44 0.0030 3.97 0.0043
-Gaussian multi-hypothesis 1.64 0.0310 2.41 0.0117 2.61 0.0036 3.71 0.0040
+(Abl.) NDT Scan Matching [1] 12.61 2.42 0.0345 7.80 0.0408 5.61 0.0065 7.84 0.0100
+(Abl.) GICP Scan Matching [2] 12.33 2.93 0.0428 2.97 0.0111 2.99 0.0038 3.50 0.0042
+(Abl.) VGICP Scan Matching [15] 6.79 2.43 0.0362 2.56 0.0136 2.73 0.0036 3.84 0.0047
+(Abl.) Gaussian Single Hypothesis 5.29 2.15 0.0318 2.46 0.0116 2.44 0.0030 3.97 0.0043
+Gaussian Multi-Hypothesis 19.17 1.64 0.0310 2.41 0.0117 2.61 0.0036 3.71 0.0040
 Table II: Results on Snail-Radar Dataset (Oculii Eagle). Best results in bold, second best results underlined.
@@
 (%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m)
-(Ablation) Our EKF with NDT [2] 4.13 0.5286 2.04 0.0519 3.95 0.0653 2.86 0.0786 3.15 0.0771
-(Ablation) Our EKF with GICP [21] 3.77 0.5118 2.93 0.0629 4.05 0.0629 2.77 0.1136 4.58 0.1572
-(Ablation) Our EKF with VGICP [12] 3.68 0.5141 3.06 0.0610 3.17 0.0551 2.53 0.1167 4.44 0.1489
-(Ablation) Gaussian single hypothesis 3.36 0.5204 2.85 0.0689 2.84 0.0537 2.36 0.0927 4.00 0.1457
+(Abl.) NDT Scan Matching [1] 4.13 0.5286 2.04 0.0519 3.95 0.0653 2.86 0.0786 3.15 0.0771
+(Abl.) GICP Scan Matching [2] 3.77 0.5118 2.93 0.0629 4.05 0.0629 2.77 0.1136 4.58 0.1572
+(Abl.) VGICP Scan Matching [15] 3.68 0.5141 3.06 0.0610 3.17 0.0551 2.53 0.1167 4.44 0.1489
+(Abl.) Gaussian Single Hypothesis 3.36 0.5204 2.85 0.0689 2.84 0.0537 2.36 0.0927 4.00 0.1457
 Gaussian Multi-Hypothesis 3.33 0.5147 2.45 0.0543 2.89 0.0498 2.55 0.1093 4.32 0.1569
@@
 (%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m)
-(Ablation) Our EKF with GICP [21] 1.69 0.3049 3.36 0.0507 3.29 0.0487 2.42 0.0761 2.26 0.0796
-(Ablation) Our EKF with VGICP [12] 1.95 0.2798 2.57 0.0494 3.10 0.0491 2.81 0.0777 2.12 0.0836
-(Ablation) Gaussian single hypothesis 1.34 0.3727 2.50 0.0487 2.77 0.0475 2.47 0.0826 2.37 0.0740
+(Abl.) GICP Scan Matching [2] 1.69 0.3049 3.36 0.0507 3.29 0.0487 2.42 0.0761 2.26 0.0796
+(Abl.) VGICP Scan Matching [15] 1.95 0.2798 2.57 0.0494 3.10 0.0491 2.81 0.0777 2.12 0.0836
+(Abl.) Gaussian Single Hypothesis 1.34 0.3727 2.50 0.0487 2.77 0.0475 2.47 0.0826 2.37 0.0740
 Gaussian Multi-Hypothesis 1.60 0.3573 4.02 0.0756 6.80 0.1000 2.93 0.0894 2.25 0.0746
@@
 VI Experimental results
-In order to evaluate the performance of our Gaussian modeling and scan matching method, we use existing public 4D radar datasets.
+We use existing public 4D radar datasets for evaluation.
 In particular, we selected the NTU4DRadLM [8] and Snail-Radar [31] datasets, because together they cover multiple radar types (Oculii Eagle and ARS548), multiple robot platform types (handcart and car), and adverse conditions.
 We evaluate relative translation/rotation errors using the modern and flexible evo evaluation package [32], while preserving existing practice by using a custom evaluation script that follows the behavior of older packages such as rpg_trajectory_evaluation as closely as possible.
-In the case of Snail-Radar sequences, we report rotation errors in mrad/m instead of º/m because the errors are much smaller.
-Besides our full method, we evaluate four ablations within our EKF framework: one without multi-hypothesis registration, and three replacing our Gaussian-based algorithms with NDT, GICP and VGICP.
-Since these are all single-hypothesis methods, we use the current nominal state of the EKF directly as the initial estimate (instead of sampling a random particle swarm).
+In the case of Snail-Radar sequences, we report rotation errors in mrad/m instead of º/m due to much smaller errors.
+Besides our Gaussian methods, we evaluate four ablations within our RIO framework: one without multi-hypothesis registration, and three using common established scan matching algorithms (NDT, GICP and VGICP) directly with radar point clouds.
+We use the current nominal state of the EKF directly as the initial estimate for single-hypothesis scan matching.
 The NDT version uses the implementation provided by the well known PCL (Point Cloud Library), while the GICP and VGICP versions use small_gicp [33] as the underlying implementation.
@@
 In both cases we sample a swarm of 8 particles during scan matching.
-The source code of our odometry solution can be accessed online on GitHub222 https://github.com/robotics-upo/gaussian-rio-cpp .
+The source code of our system, which includes the Gaussian algorithms and the odometry pipeline, can be accessed online on GitHub [^2] https://github.com/robotics-upo/gaussian-rio-cpp .
 We include the script used to generate the odometry trajectories, and the script used to evaluate trajectories with evo.
@@
 VI-A Results on NTU4DRadLM
-We compare our odometry with previous published results on the NTU4DRadLM dataset [29], which includes scans captured by an Oculii Eagle radar, and a VectorNav VN-100 IMU.
+We evaluate using the NTU4DRadLM dataset [8], which includes scans captured by an Oculii Eagle radar, and a VectorNav VN-100 IMU.
+Besides the evaluation performed with our pipeline, we also include previously published results of several radar-focused odometry methods, in order to provide context on the current performance of the state of the art.
+Due to a lack of Radar-Inertial methods, we also include radar-only methods.
 These include the GICP and Fast-LIO baselines tested in the original NTU4DRadLM paper, the APDGICP algorithm tested in the follow-up 4DRadarSLAM paper [9], as well as other works such as EFEAR-4D [10] or RIV-SLAM [11].
-We include these self-reported metrics as-is, as the trajectory files needed for metric calculation are usually not made public.
-Even though there are no published results for EKF-RIO [4] on this dataset, it is one of the few methods with available source code, and so we are able to produce and include evaluation results for it.
+All works are categorized according to the sensory input used (IMU, egovelocity, scan matching).
+We include self-reported metrics as-is, as the trajectory files needed for metric calculation are usually not made public.
+The results for EKF-RIO [29] are generated by running the system directly, as there are no published results on this dataset.
 We report metrics without loop closure whenever possible, in order to focus on evaluating pure odometry.
@@
 Table I shows the comparison between all methods.
-We can observe how our Gaussian methods provide the best or second best results for all sequences except for loop3, which seems to favor GICP-based methods (both baseline and our implementation).
+Our Gaussian methods provide the best or second best results for all sequences except for loop3, which seems to favor GICP-based methods (both baseline and our implementation).
 Moreover, in 3 out of the 4 sequences the multi-hypothesis scan matching approach provides improvements in translation error, especially in cp.
+Additionally, we report the average scan matching processing time for each method running under our odometry pipeline.
+The evaluation platform is a desktop PC equipped with an Intel Core i9-9900X CPU and 128 GB of RAM.
+The single-hypothesis version of our method achieves the fastest scan matching processing times, closely followed by VGICP.
+The multi-hypothesis approach, despite simultaneously registering 8 particles at once, is only 3.62x slower than the single-hypothesis version.
+We also mention that our current RIO implementation is entirely sequential, and that Gaussian modeling at keyframes consumes up to 608.78ms, raising the average scan processing time to 18.13ms.
+However, we believe that running the modeling and odometry tasks in parallel would remove any timing inconsistencies affecting real-time usage; which we leave as future work.
 VI-B Results on Snail-Radar
@@
 In particular, we selected the five sequences recorded on 2023/12/13 because of the nighttime and rain conditions of that day under a high speed car platform.
-These sequences are all fairly long, with the longest (iaf_20231213_2) being 15 min and most being around 7 min in length.
+The longest tested sequence (iaf_20231213_2) is 15 min in length, while the others are around 7 min.
 Table II shows results using the Oculii Eagle radar.
@@
 The multi-hypothesis approach improves or matches results compared to the single hypothesis version in 3 out of 5 sequences.
-It is worth mentioning that NDT, unlike in NTU4DRadLM, seems to behave much better in Snail-Radar.
+We also note that NDT seems to behave much better in Snail-Radar than in NTU4DRadLM.
 Table III shows results using the ARS548 radar.
 This radar produces much sparser point clouds than Oculii Eagle (around 10x fewer points).
-We believe this to be caused by the radar lacking the scan enhancement algorithm used by Oculii Eagle.
-However interestingly, we can observe better metrics in general compared to Oculii Eagle.
-The sparser scans trigger a known issue in PCL’s implementation of NDT that cause the program to abnormally terminate, so unfortunately NDT is excluded from the comparison.
+We believe this to be the effect of Oculii Eagle’s scan enhancement algorithm.
+However, we can observe better metrics in general compared to Oculii Eagle.
+The sparser scans trigger a known issue in PCL’s implementation of NDT that cause a program termination, so unfortunately NDT is excluded from the comparison.
 The results once again show different methods leading in different sequences, however it is worth pointing out that 6 out of the 10 best results are now obtained by our Gaussian algorithms, as well as 3 out of the 10 second best results.
@@
 VI-C Summary of results
-We consider that the results described above validate our methodology.
+We consider that the results above validate our methodology.
 Our Gaussian modeling and scan matching algorithm produces comparable relative translation and rotation metrics to existing registration algorithms, as shown by the experiments.
 Moreover, in several sequences across several datasets and radar types our algorithm produces the best overall results.
-Our EKF-based radar-inertial framework is also validated by the experiments, which show how on a publicly evaluated dataset we are able to improve on the state of the art even when not using our Gaussian algorithms.
 On the other hand, the multi-hypothesis approach seems to need further adjustments, although it is already capable of producing better results for some sequences.
@@
 VII Conclusion
-In this paper we introduce a Gaussian-based radar-inertial odometry pipeline.
-Our methods summarize noisy 4D radar scans using freely positioned 3D Gaussians (like 3DGS [10] and unlike NDT [2]), which then support multi-hypothesis scan matching for increased robustness.
+In this paper we introduce modeling and scan matching algorithms based on 3D Gaussians, which initialize and jointly optimize all parameters of a 3D Gaussian model to fit a radar point cloud, and make use of multiple hypothesis to support a scan matching process for subsequent incoming point clouds.
 Through experiments in multiple datasets with multiple radar types, we show that these methods produce results comparable to existing registration algorithms, and even manage to outperform them in several sequences.
-Future works include improving the Gaussian models with an optimization algorithm more efficient than gradient descent, better initialization and dynamic management of Gaussians (as in 3DGS), as well as incorporating radar-specific information such as intensity values; improving the robustness of the multi-hypothesis particle system with kinematic priors, dynamically adjusting the hyperparameters of the system, improving the quality of egovelocity estimation by modeling kinematic restrictions relevant to the robot platform used, further refining the pose estimates by means of Pose Graph Optimization (PGO), incorporating a loop closure module, and creating a global Gaussian map of the environment for the purposes of full SLAM.
+At the same time, our methods offer higher flexibility/expandability and faster scan matching compared to existing approaches based on individual points and voxelization [2, 1, 15, 16].
+Future works include improving the run-time efficiency of the Gaussian modeling (replacing gradient descent, and running in parallel to the odometry pipeline), better initialization and dynamic management of Gaussians (as in 3DGS), as well as incorporating radar-specific information such as RCS (Radar Cross Section) values, and improving the robustness of the multi-hypothesis particle system with kinematic priors.
+Another line of work is further development of the RIO pipeline, using techniques such as Pose Graph Optimization (PGO), loop closure, and creating a global Gaussian map of the environment for the purposes of full SLAM.
 Acknowledgment

### v4

@@
 However, the noisy and sparse nature of the returned scan data proves to be a challenging obstacle for existing registration algorithms, especially those originally intended for more accurate sensors such as LiDAR.
-Following the success of 3D Gaussian Splatting for vision, in this paper we propose a summarized representation for radar scenes based on global simultaneous optimization of 3D Gaussians as opposed to voxel-based approaches, and leveraging its inherent probability distribution function for registration.
-Moreover, we propose tackling the problem of radar noise by optimizing multiple scan matching hypotheses in order to further increase the robustness of the system against local optima of the function.
+Following the success of 3D Gaussian Splatting for vision, in this paper we propose a summarized representation for radar scenes based on global simultaneous optimization of 3D Gaussians as opposed to voxel-based approaches, and leveraging its inherent Probability Density Function (PDF) for registration.
+Moreover, we propose tackling the problem of radar noise entirely within the scan matching process by optimizing multiple registration hypotheses for better protection against local optima of the PDF.
 Finally, following existing practice we implement an Extended Kalman Filter-based Radar-Inertial Odometry pipeline in order to evaluate the effectiveness of our system.
@@
 However, despite working well for LiDAR [3, 4, 5], radar point clouds are considerably more challenging to register because they are sparser, noisier, and have a narrower field of view; and, as a result, these algorithms often struggle with the geometry of certain scenes.
-Figure 1: Gaussian model of an example point cloud, containing 150 Gaussians. The surface of each spheroid represents the $2\sigma$ front of each Gaussian (i.e. Mahalanobis distance equal to two).
-Figure 2: Overview of the Radar-Inertial Odometry pipeline including our proposed Gaussian modeling and scan matching processes
+Figure 1: Overview of the Radar-Inertial Odometry pipeline including our proposed Gaussian modeling and scan matching processes.
+Bold font indicates input/output, the highlighted area indicates our contribution
 We introduce a new method for modeling and registering 4D radar point clouds using 3D Gaussians.
-Following the success of 3D Gaussian Splatting (3DGS) for vision [6], we propose the use of a collection of 3D Gaussians for modeling the geometry of a scene (see Fig.
-1).
+Following the success of 3D Gaussian Splatting (3DGS) for vision [6], we propose the use of a collection of 3D Gaussians for modeling the geometry of a scene.
 As in [6] (and in contrast to NDT [1]), we simultaneously optimize all parameters of the model, including the position, scaling, rotation and correspondences between points and Gaussians.
@@
 Finally, we integrate our algorithms into a Radar-Inertial Odometry (RIO) pipeline using an Extended Kalman Filter (EKF) (see Fig.
-2), which we then use to evaluate our Gaussian methods against several established registration algorithms on multiple datasets and radar types.
+1), which we then use to evaluate our Gaussian methods against several established registration algorithms on multiple datasets and radar types.
 This paper is organized as follows: Section II reviews existing related literature, Sections III and IV describe the proposed Gaussian modeling and scan matching components respectively, Section V describes the EKF-based RIO pipeline used in the experiments; Section VI shows experimental results, and Section VII concludes the paper proposing future work.
@@
 The Normal Distributions Transform (NDT) algorithm [1] subdivides a point cloud using a voxel grid and computes normal distributions for each cell that contains points.
-Scan matching is performed by maximizing the resulting probability distribution function with respect to an incoming scan.
+Scan matching is performed by maximizing the resulting Probability Density Function (PDF) with respect to an incoming scan.
 An important limitation of this algorithm is that the grid structure itself imposes restrictions on the geometry, and the resulting model may not be fine grained enough to accurately model the geometry of the environment.
@@
 III Gaussian modeling
-Given a point cloud $\mathcal{P}$ contaning $M$ points $\{\mathbf{p}_{i}|i=0,1,\ldots,M\}$ with $\mathbf{p}_{i}\in\mathbb{R}^{3}$, our goal is creating a summarized representation of its geometry for use with downstream tasks such as mapping or localization.
+Given a point cloud $\mathcal{P}$ containing $M$ points $\{\mathbf{p}_{i}|i=0,1,\ldots,M\}$ with $\mathbf{p}_{i}\in\mathbb{R}^{3}$, our goal is creating a summarized representation of its geometry for use with downstream tasks such as mapping or localization.
 We model the geometry of a given radar scene using normal distributions, as in [1, 6, 15, 16].
-Figure 1 shows an example of a Gaussian model.
 III-A Parametrization
@@
   • Center point $\boldsymbol{\mu}_{j}\in\mathbb{R}^{3}$, indicating the mean of the distribution.
-  • Scaling vector $\mathbf{s}_{j}\in\mathbb{R}^{3}$, indicating the natural logarithms of the square roots of the eigenvalues of the covariance matrix.
+  • Scaling vector $\mathbf{s}_{j}\in\mathbb{R}^{3}$, indicating the standard deviations in each dimension.
 We impose a minimum size restriction so that Gaussians do not become infinitely small (and Mahalanobis distances infinitely large).
-  • Rotation quaternion $\mathbf{q}_{j}\in\mathbb{H}$, compactly encoding the eigenvectors of the covariance matrix (see below); where $\mathbb{H}$ is the field of quaternions.
+  • Rotation quaternion $\mathbf{q}_{j}\in\mathbb{H}$, compactly encoding the orientation of the Gaussian.
 In order to ensure that the quaternions remain unitary during the optimization, we normalize them prior to usage.
-Unlike in Gaussian Mixture Models (GMMs), we do not define $\pi_{j}$ weight parameters for the prior probability of a point being generated by a given Gaussian.
+Unlike Gaussian Mixture Models (GMMs), we do not define $\pi_{j}$ weight parameters for the prior probability of a point being generated by a given Gaussian.
 This is because point cloud data does not encode a complete population of points – it merely represents a sample of said population, biased by factors such as the physical geometric properties of the sensor.
 For instance, sampling density is inversely proportional to distance from the sensor.
-This means we cannot model these prior probabilities, and as such we only consider posterior probabilities after a point is assigned to a Gaussian (through other non-probabilistic means, such as measuring distances).
+This means we cannot model these prior probabilities, and as such we only consider posterior probabilities after a point is assigned to a Gaussian (using e.g.
+the shortest Euclidean distance).
 The scaling and rotation parameters succinctly describe the covariance matrix of a normal distribution [26, 27].
-To see why this is possible, we first derive a rotation matrix $\mathbf{R}_{j}=\mathbf{R}\{\mathbf{q}_{j}\}$, and a scaling matrix $\mathbf{S}_{j}=\mathrm{diag}(\exp_{\circ}(\mathbf{s}_{j}))$, where $\exp_{\circ}(\mathbf{x})$ is element-wise natural exponentiation and $\mathbf{R}\{\mathbf{q}\}$ is the quaternion-to-rotation-matrix conversion operator.
-We then define a transformation matrix $\mathbf{M}_{j}=\mathbf{R}_{j}\mathbf{S}_{j}$ that maps a mean-centered point into the Gaussian’s correlation space, and finally derive the covariance matrix $\boldsymbol{\Sigma}_{j}=\mathbf{M}_{j}\mathbf{M}_{j}^{T}$ of the distribution.
+We first derive a rotation matrix $\mathbf{R}_{j}=\mathbf{R}\{\mathbf{q}_{j}\}$, and a scaling matrix $\mathbf{S}_{j}=\mathrm{diag}(\exp_{\circ}(\mathbf{s}_{j}))$, where $\exp_{\circ}(\mathbf{x})$ is element-wise natural exponentiation and $\mathbf{R}\{\mathbf{q}\}$ is the rotation matrix derived from a given quaternion.
+We then define a matrix $\mathbf{M}_{j}=\mathbf{R}_{j}\mathbf{S}_{j}$ that maps a mean-centered point into the Gaussian’s correlation space, and finally derive the covariance matrix $\boldsymbol{\Sigma}_{j}=\mathbf{M}_{j}\mathbf{M}_{j}^{T}$ of the distribution.
 III-B Initialization
@@
 We propose choosing $N$ proportionally based on the number of points, so that richer scans can be represented with more Gaussians.
+The proportion should also consider the sparsity of the point clouds generated by the radar used.
 The model $\mathcal{G}$ is then initialized with all $\mathbf{s}_{j}$ set to 0 (indicating unit scale), and all $\mathbf{q}_{j}$ set to the unit quaternion.
@@
   • The parameters of $\mathcal{G}$ (in other words, all $\theta_{j}$) are updated using the gradient of the loss function.
-We define the following loss function affecting a given $\theta_{j}$, derived from the log PDF (Probability Density Function) of the multivariate normal distribution:
+We define the following loss function for each $\theta_{j}$, derived from the log PDF of the multivariate normal distribution:
 $\mathcal{L}_{j}=\frac{1}{2|G_{j}|}\sum_{\mathbf{p}_{i}\in G_{j}}\hat{\mathbf{p}}_{i}^{T}\hat{\mathbf{p}}_{i}+\sum_{k}^{3}\mathbf{s}_{j_{k}}.$ (1)
@@
 The algorithm is based on the simultaneous optimization of multiple hypotheses centered around the current predicted position of the robot, in order to increase the robustness against local optima.
+At the end of the optimization, the hypotheses are judged against a scoring function, and the best hypothesis is selected as the output of the scan matching algorithm while the other hypotheses are discarded.
+Note that this is unlike a classical particle filter, which does persist the hypotheses throughout time.
 IV-A Definitions
@@
 The criteria are then simply that any of the following must hold:
-  • For the translational component: $\mathbf{t}||_{2}\geq d_{\max}$, where $d_{\max}$ is the maximum distance threshold.
-  • For the rotational component: $2\cos^{-1}|q_{w}|\geq\alpha_{\max}$, where $\alpha_{\max}$ is the maximum rotation angle threshold.
+  • For the translational component: $\mathbf{t}||_{2}\geq{t}_{\max}$, where ${t}_{\max}$ is the maximum translation threshold.
+  • For the rotational component: $2\cos^{-1}|q_{w}|\geq\alpha_{\max}$, where $\alpha_{\max}$ is the maximum rotation threshold.
 As a fallback, if no scans have returned a successful match during the past given time interval $t$, a keyframe is automatically generated.
@@
 where $\mathbf{G}$ is the error reset matrix, and it is a sparse matrix equal to $\mathbf{I}$ except for every $\delta\theta\,\delta\theta$ subblock.
-Finally, in order to prevent especially noisy observations from causing instability in the filter, we perform Mahalanobis outlier rejection, using a threshold calculated from a given percentile of the $\chi^{2}$ distribution, like [29].
+Finally, in order to prevent especially noisy observations from causing instability in the filter, we follow [29] and perform Mahalanobis outlier rejection, using a threshold calculated from a given percentile of the $\chi^{2}$ distribution.
 V-D Radar egovelocity observation
@@
 We use the RANSAC-LSQ algorithm [29] to both filter out outlier (dynamic) points, and estimate the current radar-frame velocity of the robot, including a covariance matrix.
-The filtered point cloud is also fed to the Gaussian modeling and scan matching algorithm in the following subsection.
+The filtered point cloud is then fed to the Gaussian modeling and scan matching algorithm in the following subsection.
 The observation model used to perform the EKF update is as follows:
@@
 In other words, given a keyframe pose $\xi^{k}$ we can observe the change in translation and rotation that has occurred since said keyframe.
-We initialize the pose particle swarm with a distribution $\mathbf{P}\thicksim\mathcal{N}(\xi^{x},\boldsymbol{\Sigma})$ where $\xi^{x}$ is the relative robot pose formed from the current nominal state of the EKF: $\xi^{x}=\{\mathbf{p},\mathbf{q}\}\ominus\xi^{k}$, and $\boldsymbol{\Sigma}$ is a covariance matrix indicating the desired dispersion of the particles for each of the 6 degrees of freedom (X/Y/Z/roll/pitch/yaw).
-Afterwards, we apply the scan matching algorithm and obtain an optimized observation pose $\xi^{y}$, from which a residual pose can be obtained: $\xi^{r}=\xi^{y}\ominus\xi^{x}=\{\delta\mathbf{p},\delta\mathbf{q}\}$.
+We randomly initialize the pose particle swarm with a distribution $\mathbf{P}\thicksim\mathcal{N}(\xi^{x},\boldsymbol{\Sigma})$ where $\xi^{x}$ is the relative robot pose formed from the current nominal state of the EKF: $\xi^{x}=\{\mathbf{p},\mathbf{q}\}\ominus\xi^{k}$, and $\boldsymbol{\Sigma}$ is a covariance matrix indicating the desired dispersion of the particles for each of the 6 degrees of freedom (X/Y/Z/roll/pitch/yaw).
+Afterwards, we apply the scan matching algorithm and obtain a single optimized observation pose $\xi^{y}$, from which a residual pose can be obtained: $\xi^{r}=\xi^{y}\ominus\xi^{x}=\{\delta\mathbf{p},\delta\mathbf{q}\}$.
 Since the attitude error quaternion $\delta\mathbf{q}$ can be approximated as $\delta\mathbf{q}=\exp\left(\frac{1}{2}\delta\boldsymbol{\theta}\right)\approx[1\,\frac{1}{2}\delta\boldsymbol{\theta}^{T}]^{T}$ for very small errors, the attitude error vector can be approximated as $\delta\boldsymbol{\theta}\approx\frac{2}{\delta q_{w}}\delta\mathbf{q}_{v}$.
@@
 The non-zero subblocks of the observation matrix $\mathbf{H}$ can be defined as such: $\mathbf{H}_{\delta p\,p}=\mathbf{H}_{\delta\theta\,\delta\theta}=\mathbf{C}_{w}^{k}$, where $k$ corresponds to the keyframe’s reference frame.
-We are currently using a fixed observation covariance matrix $\mathbf{R}$, although we believe it should be possible to leverage the optimized particle swarm returned by the scan matching algorithm to estimate this matrix.
-Table I: Results on NTU4DRadLM Dataset (Oculii Eagle). Best results in bold, second best results underlined.
-Methods marked with * are self-reported metrics. EV: Egovelocity, SM: Scan Matching, Abl.: Ablation
-cp nyl loop2 loop3
-Usage Time $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
-IMU EV SM (ms) (%) (º/m) (%) (º/m) (%) (º/m) (%) (º/m)
+We currently use a fixed observation covariance matrix $\mathbf{R}$, although we believe it is possible to leverage the optimized particle swarm returned by the scan matching algorithm to estimate this matrix.
+Table I: Results on NTU4DRadLM Dataset (Oculii Eagle).
+Best results in bold, second best results underlined.
+Methods marked with * are self-reported metrics.
+EV: Egovelocity, SM: Scan Matching, Abl.: Ablation.
+Relative translation errors ($t_{\mathrm{rel}}$) are expressed in %, relative rotation errors ($r_{\mathrm{rel}}$) in °/m, and time in ms
+Radar/Method Inputs cp nyl loop2 loop3
+Oculii Eagle IMU EV SM Time $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
 NTU4DRadLM Baseline: GICP* [8] 4.13 0.0552 4.62 0.0184 4.84 0.0060 3.22 0.0060
@@
 Gaussian Multi-Hypothesis 19.17 1.64 0.0310 2.41 0.0117 2.61 0.0036 3.71 0.0040
-Table II: Results on Snail-Radar Dataset (Oculii Eagle). Best results in bold, second best results underlined.
-st_20231213_1 iaf_20231213_2 iaf_20231213_3 if_20231213_4 if_20231213_5
-$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
-(%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m)
+Table II: Results on Snail-Radar Dataset (Oculii Eagle and ARS548).
+Best results in bold, second best results underlined.
+Relative translation errors ($t_{\mathrm{rel}}$) are expressed in % and relative rotation errors ($r_{\mathrm{rel}}$) in mrad/m
+Radar/Method st_20231213_1 iaf_20231213_2 iaf_20231213_3 if_20231213_4 if_20231213_5
+Oculii Eagle $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
 (Abl.) NDT Scan Matching [1] 4.13 0.5286 2.04 0.0519 3.95 0.0653 2.86 0.0786 3.15 0.0771
@@
 Gaussian Multi-Hypothesis 3.33 0.5147 2.45 0.0543 2.89 0.0498 2.55 0.1093 4.32 0.1569
-Table III: Results on Snail-Radar Dataset (ARS548). Best results in bold, second best results underlined.
-st_20231213_1 iaf_20231213_2 iaf_20231213_3 if_20231213_4 if_20231213_5
-$t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$ $t_{rel}$ $r_{rel}$
-(%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m) (%) (mrad/m)
+ARS548 $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
 (Abl.) GICP Scan Matching [2] 1.69 0.3049 3.36 0.0507 3.29 0.0487 2.42 0.0761 2.26 0.0796
@@
 Gaussian Multi-Hypothesis 1.60 0.3573 4.02 0.0756 6.80 0.1000 2.93 0.0894 2.25 0.0746
-Furthermore, we constrain the observation model described above by removing degrees of freedom that are deemed noisy in the scan matching due to the properties of the sensor.
+Furthermore, we follow [31] and constrain the observation model described above by removing degrees of freedom that are deemed noisy in the scan matching due to the properties of the sensor.
 For example, radars have lower resolution in the Z axis, which degrades the usefulness of Z/roll/pitch information returned by scan matching algorithms.
@@
 In the case of Oculii Eagle scans, we set a target of 20 points per Gaussian; while for ARS548 scans, due to the much lower number of returned points, we set a target of 8 points per Gaussian.
-In both cases we sample a swarm of 8 particles during scan matching.
+In both cases we sample a swarm of 8 particles during scan matching, which we believe are enough for testing purposes without significantly increasing processing times.
+Other parameters include $t_{\mathrm{max}}$ = 15m and $\alpha_{\mathrm{max}}$ = 5° for keyframing, $d_{\mathrm{max}}$ = 4 for scan matching, and 0.5m/2° for multi-hypothesis particle dispersion.
+These parameters were chosen to suit the typical depth range/sparsity relation of a radar.
 The source code of our system, which includes the Gaussian algorithms and the odometry pipeline, can be accessed online on GitHub [^2] https://github.com/robotics-upo/gaussian-rio-cpp .
@@
 We also note that NDT seems to behave much better in Snail-Radar than in NTU4DRadLM.
-Table III shows results using the ARS548 radar.
+Table II also shows results using the ARS548 radar.
 This radar produces much sparser point clouds than Oculii Eagle (around 10x fewer points).
@@
 We hypothesize this to be related to the increased sparsity of the scans, and believe it could be solved by improving the sampling distribution of the particles.
-VI-C Summary of results
+VI-C Additional qualitative experiments
+In order to further justify our modeling approach based on globally optimized 3D Gaussians versus existing voxel-based approaches, we compare in Fig.
+2 the output of our method against NDT.
+To do this, we prepare a radar-based map of NTU4DRadLM’s cp sequence using the supplied ground truth, and model a section of said map with both NDT and our method.
+The positions and sizes of the Gaussians generated by NDT are clearly constrained and influenced by the voxel grid, while our approach generates better fitted Gaussians.
+Moreover, our method is able to preserve geometric details such as corners, and allocates a higher budget of Gaussians to areas with higher point density, resulting in overall better representation of the geometry.
+Additionally, we include a video supplement.
+In the video we show several Gaussian models of the same map created using different points-per-Gaussian ratios.
+It can be seen that, regardless of said hyperparameter, our modeling method is able to generate a representation reasonably fitting the map.
+Different number of Gaussians are allocated to areas of the map depending on the density and complexity of the area, as well as the available budget.
+Moreover, when the ratio is higher (meaning fewer Gaussians), sparse regions with few (outlier) radar points do not receive any Gaussians.
+Finally, we also include an example of the radar-inertial odometry pipeline on the same NTU4DRadLM cp sequence, compared to the ground truth trajectory.
+Figure 2: Qualitative comparison between NDT and our Gaussian modeling using NTU4DRadLM’s cp sequence.
+VI-D Summary of results
 We consider that the results above validate our methodology.

### v5

@@
 Following the success of 3D Gaussian Splatting for vision, in this paper we propose a summarized representation for radar scenes based on global simultaneous optimization of 3D Gaussians as opposed to voxel-based approaches, and leveraging its inherent Probability Density Function (PDF) for registration.
-Moreover, we propose tackling the problem of radar noise entirely within the scan matching process by optimizing multiple registration hypotheses for better protection against local optima of the PDF.
-Finally, following existing practice we implement an Extended Kalman Filter-based Radar-Inertial Odometry pipeline in order to evaluate the effectiveness of our system.
+Moreover, we propose optimizing multiple registration hypotheses for better protection against local optima of the PDF.
+We evaluate our modeling and registration system against state of the art techniques, finding that our system provides richer models and more accurate registration results.
+Finally, we evaluate the effectiveness of our system in a real Radar-Inertial Odometry task.
 Experiments using publicly available 4D radar datasets show that our Gaussian approach is comparable to existing registration algorithms, outperforming them in several sequences.
@@
 However, despite working well for LiDAR [3, 4, 5], radar point clouds are considerably more challenging to register because they are sparser, noisier, and have a narrower field of view; and, as a result, these algorithms often struggle with the geometry of certain scenes.
-Figure 1: Overview of the Radar-Inertial Odometry pipeline including our proposed Gaussian modeling and scan matching processes.
-Bold font indicates input/output, the highlighted area indicates our contribution
 We introduce a new method for modeling and registering 4D radar point clouds using 3D Gaussians.
@@
 This approach is entirely featureless, and does not require any priors other than an initial set of centroids.
-Moreover, we propose a registration step that considers multiple simultaneous hypotheses in order to add further robustness against radar noise.
-Finally, we integrate our algorithms into a Radar-Inertial Odometry (RIO) pipeline using an Extended Kalman Filter (EKF) (see Fig.
-1), which we then use to evaluate our Gaussian methods against several established registration algorithms on multiple datasets and radar types.
-This paper is organized as follows: Section II reviews existing related literature, Sections III and IV describe the proposed Gaussian modeling and scan matching components respectively, Section V describes the EKF-based RIO pipeline used in the experiments; Section VI shows experimental results, and Section VII concludes the paper proposing future work.
+Moreover, we propose a registration step that allows for multiple simultaneous hypotheses in order to add further robustness against local optimization minima.
+We evaluate our Gaussian approach against several established registration algorithms with a test suite independently evaluating the modeling and scan matching, as well as combined within a full Radar-Inertial Odometry (RIO) pipeline using an Extended Kalman Filter (EKF) on multiple datasets and radar types.
+This paper is organized as follows: Section II reviews existing related literature, Sections III and IV describe the proposed Gaussian modeling and scan matching components respectively, Section V shows experimental results, and Section VI concludes the paper proposing future work.
 II Related work
@@
 As for $\boldsymbol{\mu}_{j}$, we perform Bisecting K-Means clustering to find initial estimates for the centers of the Gaussians.
+Compared to voxel-based clustering used by NDT or VGICP [1, 15], this approach produces fewer spurious Gaussians from outlier points, as K-Means is biased towards higher density areas.
 III-C Optimization
@@
 The loss function $\mathcal{L}$ of the entire model $\mathcal{G}$ can be obtained by taking the mean of the loss values across all Gaussians.
-IV Gaussian multi-hypothesis scan matching
+IV Gaussian scan matching
 Given a Gaussian model $\mathcal{G}$ and a point cloud $\mathcal{P}$, we define a registration process that estimates the pose of the robot $\xi$ with respect to $\mathcal{G}$.
-The algorithm is based on the simultaneous optimization of multiple hypotheses centered around the current predicted position of the robot, in order to increase the robustness against local optima.
+Moreover, we extend the usual approach to support the simultaneous optimization of multiple hypotheses centered around the current predicted position of the robot, in order to increase the robustness against local optima.
 At the end of the optimization, the hypotheses are judged against a scoring function, and the best hypothesis is selected as the output of the scan matching algorithm while the other hypotheses are discarded.
@@
 We consider that the uncertainty about the true robot pose follows some probability distribution: $\xi\thicksim\mathrm{P}$.
-Given a suitable model of $\mathrm{P}$, we can randomly sample a swarm of $K$ pose particles $\hat{\xi}^{k}$, each representing a different hypothesis, and use an optimization algorithm that evaluates each particle against $\mathcal{G}$ to further refine these estimates until one of the particles reaches a suitable optimum.
-We believe exploring the solution space using a particle swarm increases robustness in challenging situations, especially when dealing with large rotations or translations.
+Given a suitable model of $\mathrm{P}$, we can sample a swarm of $K$ pose particles $\hat{\xi}^{k}$, each representing a different hypothesis, and use an optimization algorithm that evaluates each particle against $\mathcal{G}$ to further refine these estimates until one of the particles reaches a suitable optimum.
 IV-B Optimization
@@
   • We perform the usual Gauss-Newton update step for each particle $\hat{\xi}^{k}$, using the sums of squared Mahalanobis distances as the function to optimize, as in GICP.
-We also exclude points whose $d^{k}_{i}$ is greater than a given distance threshold $d_{\textrm{max}}$.
+Each point contributes to the gradient according to the following weight:
+    $w^{k}_{i}=\min\left(1,\frac{d_{\textrm{max}}}{d^{k}_{i}}\right);$ (2)
+    where $d_{\textrm{max}}$ is a given Mahalanobis distance threshold.
 The optimization stops when all particles converge to a solution (i.e.
@@
 The resulting output pose of the registration algorithm is the particle with the lowest value of the scoring function $\mathcal{L}_{k}$ defined as follows:
-$\mathcal{L}_{k}=\frac{1}{M}\sum_{i}^{M}\min(d^{k}_{i},d_{\textrm{max}}),$ (2)
+$\mathcal{L}_{k}=\frac{1}{M}\sum_{i}^{M}\min\left(d^{k}_{i},d_{\textrm{max}}\right),$ (3)
 where $M$ is the number of points in the scan. We consider a failure condition when the output particle has not converged.
-IV-C Keyframing
-A common technique used in odometry systems is keyframing.
-Certain poses of the robot are selected to become local reference points (keyframes) that are typically used for registration, pose graph optimization or (in the case of SLAM) loop closure purposes.
-In our case, we only use keyframing to select individual radar point clouds to undergo Gaussian modeling, and serve as reference for Gaussian scan matching of subsequent radar point clouds.
-The system only retains the pose and Gaussian model of the most recently generated keyframe.
-This approach, compared to trivial successive applications of the scan matching algorithm, has the benefit of reducing the accumulated error, though it also means the criteria used to select keyframe candidates needs to be carefully designed to generate an appropriate keyframe pacing.
-We use a simple set of criteria to decide when a new pose $\xi^{j}$ qualifies to become a new keyframe.
-Given the most recent keyframe $\xi^{i}$, we can derive the $SE(3)$ transformation relative to the keyframe $T=\xi^{j}\ominus\xi^{i}=\{\mathbf{t},\mathbf{q}\}$.
-The criteria are then simply that any of the following must hold:
-  • For the translational component: $\mathbf{t}||_{2}\geq{t}_{\max}$, where ${t}_{\max}$ is the maximum translation threshold.
-  • For the rotational component: $2\cos^{-1}|q_{w}|\geq\alpha_{\max}$, where $\alpha_{\max}$ is the maximum rotation threshold.
-As a fallback, if no scans have returned a successful match during the past given time interval $t$, a keyframe is automatically generated.
-V Inertial odometry with EKF
-In order to evaluate the suitability and effectiveness of the Gaussian algorithms for Radar-Inertial Odometry, it is necessary to define a pipeline for use in the experiments.
-Following well established practice [28, 29, 12], we use an Extended Kalman Filter in our odometry solution to tightly integrate all incoming sensor information: inertial data, radar egovelocity observations, and scan matching results.
-In order to simplify the formulation, we only use Error State filtering for rotational variables, as opposed to the entire state.
-We employ the right-handed North-West-Up (NWU) convention typical of robotics, and consider three reference frames: world ($w$), body/IMU ($b$) and radar ($r$).
-The rotation matrix $\mathbf{C}^{g}_{f}$ transforms a vector from frame $f$ to frame $g$, likewise $\mathbf{C}^{f}_{g}=(\mathbf{C}^{g}_{f})^{T}$.
-We decompose each $\mathbf{C}$ into an error part and a nominal quaternion part as follows: $\mathbf{C}=\exp\left([\delta\boldsymbol{\theta}]_{\times}\right)\mathbf{R}\{\mathbf{q}\}$, approximating the error part as $\mathbf{I}+[\delta\boldsymbol{\theta}]_{\times}$ for the purposes of linearization.
-The state ($\mathbf{x}$), control ($\mathbf{u}$) and noise ($\mathbf{w}$) vectors of the filter are the following:
-$\mathbf{x}=\begin{bmatrix}\mathbf{p}\\ \mathbf{v}\\ \mathbf{t}_{r}^{b}\\ \mathbf{b}_{a}\\ \mathbf{b}_{\omega}\\ \delta\boldsymbol{\theta}_{b}^{w}\\ \delta\boldsymbol{\theta}_{r}^{b}\end{bmatrix};\;\mathbf{u}=\begin{bmatrix}\mathbf{a}\\ \boldsymbol{\omega}\end{bmatrix};\;\mathbf{w}=\begin{bmatrix}\mathbf{w}_{v}\\ \mathbf{w}_{\theta}\\ \mathbf{w}_{a}\\ \mathbf{w}_{\omega}\\ \mathbf{w}_{b_{a}}\\ \mathbf{w}_{b_{\omega}}\end{bmatrix}$, (3)
-where $\mathbf{p}$ is the position, $\mathbf{v}$ is the velocity, $\mathbf{t}_{r}^{b}$ is the translation between radar and IMU, $\mathbf{b}_{a}$ and $\mathbf{b}_{\omega}$ are the accelerometer/gyroscope biases, each $\delta\boldsymbol{\theta}$ is the error part of its corresponding $\mathbf{C}$ frame expressed as a tangential rotation $\in\mathfrak{so}(3)$, and $\mathbf{a}$ and $\boldsymbol{\omega}$ are the IMU accelerometer/gyroscope readings.
-The position and velocity of the robot are expressed in the world frame, while the nominal rotations $\mathbf{q}_{b}^{w}$ and $\mathbf{q}_{r}^{b}$ are stored outside the filter state as quaternions.
-Velocity and attitude process noise are modeled by $\mathbf{w}_{v}$ and $\mathbf{w}_{\theta}$ respectively.
-We follow Kalibr’s IMU noise model [30] [^1] https://github.com/ethz-asl/kalibr/wiki/IMU-Noise-Model, and consider both additive ($\mathbf{w}_{a}$, $\mathbf{w}_{\omega}$) and random walk noise ($\mathbf{w}_{b_{a}}$, $\mathbf{w}_{b_{\omega}}$) in accelerometer/gyroscope readings.
-There also exists a vector $\mathbf{g}=-[0\;0\;9.80511]^{T}$ modeling the Earth’s gravity treated as a constant, as we do not consider long enough travel distances for changes in the plane tangent to the Earth’s surface to matter.
-V-A EKF initialization
-First, we initialize all components of the state vector $\mathbf{x}$ to 0, and the rotational variables to the unit quaternion.
-The initial radar-to-IMU frame ($\mathbf{t}_{r}^{b}$, $\mathbf{q}_{r}^{b}$) can be supplied externally.
-The robot is assumed to remain still during the first few seconds, which allows for averaging initial accelerometer/gyroscope readings and extracting an estimate of the body-frame gravity vector $\tilde{\mathbf{g}}_{b}$ (after flipping sign) and the gyroscope bias $\mathbf{b}_{\omega}$ respectively.
-An initial value for the roll/pitch components of $\mathbf{q}_{b}^{w}$ can then be derived from $\tilde{\mathbf{g}}_{b}$, while the yaw component is set to 0 since we are not incorporating a compass/magnetometer.
-In addition, we derive the initial accelerometer bias as $\mathbf{b}_{a}=\mathbf{C}_{w}^{b}\left(\mathbf{C}_{b}^{w}\tilde{\mathbf{g}}_{b}-\mathbf{g}\right)$, thus taking into account the difference in magnitude between the expected and observed gravity vectors.
-We initialize the covariance matrix mostly as 0, except for the subblocks related to $\mathbf{t_{r}^{b}}$, $\mathbf{b}_{a}$, $\mathbf{b}_{\omega}$, $\delta\boldsymbol{\theta}_{b}^{w}$ and $\delta\boldsymbol{\theta}_{r}^{b}$, whose diagonals we fill in according to supplied initial values.
-V-B EKF propagation
-We use a simplified inertial strapdown formulation that assumes movement to be linear instead of being perturbed by the current angular velocity reading of the gyroscope, due to the short time interval and small change in attitude (in effect equivalent to first order integration):
-$\mathbf{p}$ $\leftarrow\mathbf{p}+\mathbf{v}t+\frac{1}{2}(\mathbf{C}^{w}_{b}(\mathbf{a}-\mathbf{b}_{a})+\mathbf{g})t^{2},$ (4)
-$\mathbf{v}$ $\leftarrow\mathbf{v}+(\mathbf{C}^{w}_{b}(\mathbf{a}-\mathbf{b}_{a})+\mathbf{g})t,$ (5)
-$\mathbf{q}$ $\leftarrow\mathbf{q}\otimes\exp\left(\frac{1}{2}(\boldsymbol{\omega}-\mathbf{b}_{\omega})t\right).$ (6)
-The strapdown formulation is linearized in order to propagate the covariance matrix:
-$\mathbf{P}\leftarrow\mathbf{F}\mathbf{P}\mathbf{F}^{T}+\mathbf{N}\mathbf{Q}\mathbf{N}^{T},$ (7)
-where $\mathbf{F}$ is the linearized state transition matrix, $\mathbf{Q}$ is the noise covariance matrix, and $\mathbf{N}$ is the noise state transition matrix.
-These three matrices are sparsely populated: $\mathbf{F}=\mathbf{I}$ and $\mathbf{Q}=\mathbf{N}=\textbf{0}$, except for the following subblocks:
-$\mathbf{F}_{p\,v}=\mathbf{I}t;\;\mathbf{F}_{p\,b_{a}}=-\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{F}_{p\,\delta\theta_{b}^{w}}=-\frac{1}{2}[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t^{2};$ (8)
-$\mathbf{F}_{v\,b_{a}}=\mathbf{F}_{\delta\theta_{b}^{w}\,b_{\omega}}=-\mathbf{C}^{w}_{b}t;\;\mathbf{F}_{v\,\delta\theta_{b}^{w}}=-[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t;$ (9)
-$\mathbf{Q}_{v\,v}=\mathbf{I}\sigma^{2}_{v};\;\mathbf{Q}_{\theta\,\theta}=\mathbf{I}\sigma^{2}_{\theta};\mathbf{Q}_{a\,a}=\mathbf{I}\frac{\sigma^{2}_{a}}{t};\;$ (10)
-$\mathbf{Q}_{\omega\,\omega}=\mathbf{I}\frac{\sigma^{2}_{\omega}}{t};\mathbf{Q}_{b_{a}\,b_{a}}=\mathbf{I}\sigma^{2}_{b_{a}}t;\;\mathbf{Q}_{b_{\omega}\,b_{\omega}}=\mathbf{I}\sigma^{2}_{b_{\omega}}t;$ (11)
-$\mathbf{N}_{p\,a}=\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{N}_{v\,a}=\mathbf{N}_{\delta\theta_{b}^{w}\,\omega}=\mathbf{C}^{w}_{b}t;$ (12)
-$\mathbf{N}_{v\,v}=\mathbf{N}_{\delta\theta_{b}^{w}\,\theta}=\mathbf{N}_{b_{a}\,b_{a}}=\mathbf{N}_{b_{\omega}\,b_{\omega}}=\mathbf{I},$ (13)
-where $\sigma_{v}$, $\sigma_{\theta}$, $\sigma_{a}$, $\sigma_{\omega}$, $\sigma_{b_{a}}$ and $\sigma_{b_{\omega}}$ are the standard deviations of the respective components of $\mathbf{w}$.
-V-C EKF update and error state reset
-Given observation $\mathbf{y}$, residual observation vector $\mathbf{r}$, observation covariance matrix $\mathbf{R}$ and linearized observation matrix $\mathbf{H}$ (so that $\mathbf{r}\approx\mathbf{y}-\mathbf{H}\mathbf{x}$), we can proceed to apply the Kalman filter equations (in Joseph form):
-$\mathbf{K}=\mathbf{P}\mathbf{H}^{T}(\mathbf{H}\mathbf{P}\mathbf{H}^{T}+\mathbf{R})^{-1};\;\mathbf{L}=\mathbf{I}-\mathbf{K}\mathbf{H},$ (14)
-$\mathbf{x}\leftarrow\mathbf{x}+\mathbf{K}\mathbf{r};\;\mathbf{P}\leftarrow\mathbf{L}\mathbf{P}\mathbf{L}^{T}+\mathbf{K}\mathbf{R}\mathbf{K}^{T}.$ (15)
-As previously mentioned, rotations are expressed as error state in the EKF, and after an EKF update the nominal value of each $\delta\boldsymbol{\theta}$ becomes non-zero, making it necessary to incorporate the estimated error back into the state as follows:
-$\mathbf{q}\leftarrow\delta\mathbf{q}\otimes\mathbf{q};\;\delta\mathbf{q}=\exp\left(\frac{1}{2}\delta\boldsymbol{\theta}\right),$ (16)
-$\delta\boldsymbol{\theta}\leftarrow 0;\;\mathbf{P}\leftarrow\mathbf{G}\mathbf{P}\mathbf{G}^{T};\;\mathbf{G}_{\delta\theta\,\delta\theta}=\mathbf{R}\{\delta\mathbf{q}\},$ (17)
-where $\mathbf{G}$ is the error reset matrix, and it is a sparse matrix equal to $\mathbf{I}$ except for every $\delta\theta\,\delta\theta$ subblock.
-Finally, in order to prevent especially noisy observations from causing instability in the filter, we follow [29] and perform Mahalanobis outlier rejection, using a threshold calculated from a given percentile of the $\chi^{2}$ distribution.
-V-D Radar egovelocity observation
-Following previous works, we directly observe the robot’s velocity by leveraging the Doppler effect in measurements returned by the radar.
-We use the RANSAC-LSQ algorithm [29] to both filter out outlier (dynamic) points, and estimate the current radar-frame velocity of the robot, including a covariance matrix.
-The filtered point cloud is then fed to the Gaussian modeling and scan matching algorithm in the following subsection.
-The observation model used to perform the EKF update is as follows:
-$\mathbf{r}=\mathbf{y}-\mathbf{C}_{b}^{r}\left([\boldsymbol{\omega}]_{\times}\mathbf{t}_{r}^{b}+\mathbf{C}^{b}_{w}\mathbf{v}\right);$ (18)
-$\mathbf{H}_{:\,v}=\mathbf{C}_{b}^{r}\mathbf{C}_{w}^{b};\;\mathbf{H}_{:\,t_{r}^{b}}=\mathbf{C}_{b}^{r}[\boldsymbol{\omega}]_{\times};\;\mathbf{H}_{:\,b_{\omega}}=\mathbf{C}_{b}^{r}[\mathbf{t}_{r}^{b}]_{\times};$ (19)
-$\mathbf{H}_{:\,\delta\theta_{b}^{w}}=\mathbf{C}_{b}^{r}\mathbf{C}_{w}^{b}[\mathbf{v}]_{\times};\;\mathbf{H}_{:\,\delta\theta_{r}^{b}}=\mathbf{C}_{b}^{r}[[\boldsymbol{\omega}]_{\times}\mathbf{t}_{r}^{b}+\mathbf{C}^{b}_{w}\mathbf{v}]_{\times}.$ (20)
-V-E Scan matching observation
-The Gaussian modeling and scan matching algorithms allow us to observe robot poses relative to previous keyframes.
-In other words, given a keyframe pose $\xi^{k}$ we can observe the change in translation and rotation that has occurred since said keyframe.
-We randomly initialize the pose particle swarm with a distribution $\mathbf{P}\thicksim\mathcal{N}(\xi^{x},\boldsymbol{\Sigma})$ where $\xi^{x}$ is the relative robot pose formed from the current nominal state of the EKF: $\xi^{x}=\{\mathbf{p},\mathbf{q}\}\ominus\xi^{k}$, and $\boldsymbol{\Sigma}$ is a covariance matrix indicating the desired dispersion of the particles for each of the 6 degrees of freedom (X/Y/Z/roll/pitch/yaw).
-Afterwards, we apply the scan matching algorithm and obtain a single optimized observation pose $\xi^{y}$, from which a residual pose can be obtained: $\xi^{r}=\xi^{y}\ominus\xi^{x}=\{\delta\mathbf{p},\delta\mathbf{q}\}$.
-Since the attitude error quaternion $\delta\mathbf{q}$ can be approximated as $\delta\mathbf{q}=\exp\left(\frac{1}{2}\delta\boldsymbol{\theta}\right)\approx[1\,\frac{1}{2}\delta\boldsymbol{\theta}^{T}]^{T}$ for very small errors, the attitude error vector can be approximated as $\delta\boldsymbol{\theta}\approx\frac{2}{\delta q_{w}}\delta\mathbf{q}_{v}$.
-This allows us to construct the residual vector $\mathbf{r}=[\delta\mathbf{p}^{T}\,\delta\boldsymbol{\theta}^{T}]^{T}$ to be fed to the EKF update equation – note that this vector contains X/Y/Z/roll/pitch/yaw components (in this order).
-The non-zero subblocks of the observation matrix $\mathbf{H}$ can be defined as such: $\mathbf{H}_{\delta p\,p}=\mathbf{H}_{\delta\theta\,\delta\theta}=\mathbf{C}_{w}^{k}$, where $k$ corresponds to the keyframe’s reference frame.
-We currently use a fixed observation covariance matrix $\mathbf{R}$, although we believe it is possible to leverage the optimized particle swarm returned by the scan matching algorithm to estimate this matrix.
+V Experimental results
+We use existing public 4D radar datasets for evaluation.
+In particular, we selected the NTU4DRadLM [8] and Snail-Radar [28] datasets, because together they cover multiple radar types (Oculii Eagle and ARS548), multiple robot platform types (handcart and car), and adverse conditions.
+In particular, NTU4DRadLM includes scans captured by an Oculii Eagle radar, and a VectorNav VN-100 IMU; while Snail-Radar uses both Oculii Eagle and ARS548 radars together with an Xsens MTi-3-DK IMU, and includes sequences recorded in harsher environment conditions.
+We compare our Gaussian approach to three common established scan matching algorithms (NDT, GICP and VGICP).
+NDT uses the implementation provided by the well known PCL (Point Cloud Library), while GICP and VGICP use small_gicp [29] as the underlying implementation.
+In the case of Oculii Eagle scans, we set a target of 16 points per Gaussian; while for ARS548 scans, due to the much lower number of returned points, we set a target of 8 points per Gaussian.
+Moreover, these sparser scans trigger a known issue in PCL’s implementation of NDT that cause a program termination, so unfortunately NDT is excluded from experiments involving ARS548.
+Regarding our scan matching method, we provide results with 1, 8 and 128 particles wherever applicable.
+Other parameters include $d_{\mathrm{max}}$ = 4 for scan matching, and 5m/5° for particle dispersion.
+These parameters were chosen to suit the typical depth range/sparsity relation of a radar.
+V-A Modeling experiments
+In order to justify our modeling approach based on globally optimized 3D Gaussians versus existing voxel-based approaches, we compare in Fig.
+1 the output of our method against NDT.
+To do this, we prepare a radar-based map of NTU4DRadLM’s cp sequence using the supplied ground truth, and model a section of said map with both NDT and our method.
+The positions and sizes of the Gaussians generated by NDT are clearly constrained and influenced by the voxel grid, while our approach generates better fitted Gaussians.
+Moreover, our method is able to preserve geometric details such as corners, and allocates a higher budget of Gaussians to areas with higher point density, resulting in overall better representation of the geometry.
+Additionally, we include a video supplement.
+In the video we show several Gaussian models of the same map created using different points-per-Gaussian ratios.
+It can be seen that, regardless of said hyperparameter, our modeling method is able to generate a representation reasonably fitting the map.
+Different number of Gaussians are allocated to areas of the map depending on the density and complexity of the area, as well as the available budget.
+Moreover, when the ratio is higher (meaning fewer Gaussians), sparse regions with few (outlier) radar points do not receive any Gaussians, which we attribute to our use of K-Means clustering for initialization.
+Finally, we also include an example of the radar-inertial odometry pipeline on the same NTU4DRadLM cp sequence, compared to the ground truth trajectory.
+Figure 1: Qualitative comparison between NDT and our Gaussian modeling using NTU4DRadLM’s cp sequence.
+Figure 2: Results of the sensibility study, showing failure rate (%), translation error (m) and rotation error (°).
+Lower is better.
+Green dots represent mean errors, with the numerical value shown below.
 Table I: Results on NTU4DRadLM Dataset (Oculii Eagle).
 Best results in bold, second best results underlined.
+Lower is better.
 Methods marked with * are self-reported metrics.
-EV: Egovelocity, SM: Scan Matching, Abl.: Ablation.
+EV: Egovelocity, SM: Scan Matching.
 Relative translation errors ($t_{\mathrm{rel}}$) are expressed in %, relative rotation errors ($r_{\mathrm{rel}}$) in °/m, and time in ms
@@
 Oculii Eagle IMU EV SM Time $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
-NTU4DRadLM Baseline: GICP* [8] 4.13 0.0552 4.62 0.0184 4.84 0.0060 3.22 0.0060
-NTU4DRadLM Baseline: Fast-LIO* [8] 2.94 0.0468 3.80 0.0208 7.16 0.0057 4.55 0.0064
-4DRadarSLAM: APDGICP* [9] 3.56 0.0369 3.55 0.0171 6.09 0.0082 4.09 0.0097
+NTU4DRadLM baseline [8]: GICP* 4.13 0.0552 4.62 0.0184 4.84 0.0060 3.22 0.0060
+NTU4DRadLM baseline [8]: Fast-LIO* 2.94 0.0468 3.80 0.0208 7.16 0.0057 4.55 0.0064
+4DRadarSLAM [9]: APDGICP* 3.56 0.0369 3.55 0.0171 6.09 0.0082 4.09 0.0097
 EFEAR-4D* [10] 5.09 0.0125 8.93 0.0166 37.27 0.0152 37.01 0.0175
@@
 EKF-RIO [30] 14.76 0.3955 35.89 0.2464 44.31 0.0349 50.23 0.0910
-(Abl.) NDT Scan Matching [1] 12.61 2.42 0.0345 7.80 0.0408 5.61 0.0065 7.84 0.0100
-(Abl.) GICP Scan Matching [2] 12.33 2.93 0.0428 2.97 0.0111 2.99 0.0038 3.50 0.0042
-(Abl.) VGICP Scan Matching [15] 6.79 2.43 0.0362 2.56 0.0136 2.73 0.0036 3.84 0.0047
-(Abl.) Gaussian Single Hypothesis 5.29 2.15 0.0318 2.46 0.0116 2.44 0.0030 3.97 0.0043
-Gaussian Multi-Hypothesis 19.17 1.64 0.0310 2.41 0.0117 2.61 0.0036 3.71 0.0040
+NDT baseline [1] 12.61 2.42 0.0345 7.80 0.0408 5.61 0.0065 7.84 0.0100
+GICP baseline [2] 12.33 2.93 0.0428 2.97 0.0111 2.99 0.0038 3.50 0.0042
+VGICP baseline [15] 6.79 2.43 0.0362 2.56 0.0136 2.73 0.0036 3.84 0.0047
+Gaussian (ours) ($K=1$) 5.29 2.26 0.0339 2.52 0.0117 2.26 0.0034 3.82 0.0043
+Gaussian (ours) ($K=8$) 19.17 2.06 0.0266 2.71 0.0132 2.62 0.0039 3.41 0.0040
+Gaussian (ours) ($K=128$) 433.52 2.31 0.0339 2.89 0.0138 2.67 0.0042 3.43 0.0039
 Table II: Results on Snail-Radar Dataset (Oculii Eagle and ARS548).
+Lower is better.
 Best results in bold, second best results underlined.
@@
 Oculii Eagle $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
-(Abl.) NDT Scan Matching [1] 4.13 0.5286 2.04 0.0519 3.95 0.0653 2.86 0.0786 3.15 0.0771
-(Abl.) GICP Scan Matching [2] 3.77 0.5118 2.93 0.0629 4.05 0.0629 2.77 0.1136 4.58 0.1572
-(Abl.) VGICP Scan Matching [15] 3.68 0.5141 3.06 0.0610 3.17 0.0551 2.53 0.1167 4.44 0.1489
-(Abl.) Gaussian Single Hypothesis 3.36 0.5204 2.85 0.0689 2.84 0.0537 2.36 0.0927 4.00 0.1457
-Gaussian Multi-Hypothesis 3.33 0.5147 2.45 0.0543 2.89 0.0498 2.55 0.1093 4.32 0.1569
+NDT baseline [1] 4.13 0.5286 2.04 0.0519 3.95 0.0653 2.86 0.0786 3.15 0.0771
+GICP baseline [2] 3.77 0.5118 2.93 0.0629 4.05 0.0629 2.77 0.1136 4.58 0.1572
+VGICP baseline [15] 3.68 0.5141 3.06 0.0610 3.17 0.0551 2.53 0.1167 4.44 0.1489
+Gaussian (ours) ($K=1$) 3.48 0.5112 2.97 0.0658 2.79 0.0537 2.48 0.1233 4.60 0.1564
+Gaussian (ours) ($K=8$) 3.44 0.5165 4.15 0.0912 3.78 0.0676 2.33 0.0984 3.55 0.1202
+Gaussian (ours) ($K=128$) 3.55 0.5507 3.43 0.0833 2.74 0.0480 2.34 0.1050 3.89 0.1227
 ARS548 $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
-(Abl.) GICP Scan Matching [2] 1.69 0.3049 3.36 0.0507 3.29 0.0487 2.42 0.0761 2.26 0.0796
-(Abl.) VGICP Scan Matching [15] 1.95 0.2798 2.57 0.0494 3.10 0.0491 2.81 0.0777 2.12 0.0836
-(Abl.) Gaussian Single Hypothesis 1.34 0.3727 2.50 0.0487 2.77 0.0475 2.47 0.0826 2.37 0.0740
-Gaussian Multi-Hypothesis 1.60 0.3573 4.02 0.0756 6.80 0.1000 2.93 0.0894 2.25 0.0746
-Furthermore, we follow [31] and constrain the observation model described above by removing degrees of freedom that are deemed noisy in the scan matching due to the properties of the sensor.
-For example, radars have lower resolution in the Z axis, which degrades the usefulness of Z/roll/pitch information returned by scan matching algorithms.
-In order to ignore the unreliable degrees of freedom, we transform the residual vector, observation matrix and covariance matrix by removing the rows and/or columns corresponding to Z/roll/pitch prior to applying the EKF update.
-VI Experimental results
-We use existing public 4D radar datasets for evaluation.
-In particular, we selected the NTU4DRadLM [8] and Snail-Radar [32] datasets, because together they cover multiple radar types (Oculii Eagle and ARS548), multiple robot platform types (handcart and car), and adverse conditions.
+GICP baseline [2] 1.69 0.3049 3.36 0.0507 3.29 0.0487 2.42 0.0761 2.26 0.0796
+VGICP baseline [15] 1.95 0.2798 2.57 0.0494 3.10 0.0491 2.81 0.0777 2.12 0.0836
+Gaussian (ours) ($K=1$) 1.79 0.3561 3.37 0.0529 2.67 0.0380 2.37 0.0683 3.09 0.0902
+Gaussian (ours) ($K=8$) 1.79 0.3561 3.56 0.0572 3.24 0.0424 2.14 0.0733 2.97 0.0883
+Gaussian (ours) ($K=128$) 2.04 0.3580 3.67 0.0608 2.55 0.0420 2.29 0.0697 2.61 0.0986
+Figure 3: Select trajectories of sequences from NTU4DRadLM and Snail-Radar, showing both the ground truth and various different scan matching algorithms used within our odometry pipeline.
+Table III: Results of the ablation study on the modeling approach, replacing global/jointly optimized Gaussian model with local/voxelized VGICP model in our method.
+Lower is better.
+Best results in bold.
+Relative translation errors ($t_{\mathrm{rel}}$) expressed in %, relative rotation errors ($r_{\mathrm{rel}}$) in °/m (NTU4DRadLM) or mrad/m (Snail-Radar).
+Radar/Sequence VGICP model Gaussian model
+NTU4DRadLM
+Oculii Eagle $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
+cp 4.30 0.0651 2.26 0.0339
+nyl 6.66 0.0333 2.52 0.0117
+loop2 5.81 0.0067 2.26 0.0034
+loop3 6.20 0.0081 3.82 0.0043
+Avg. 5.74 0.0283 2.72 0.0133
+Snail-Radar
+Oculii Eagle $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
+st_20231213_1 3.87 0.4654 3.48 0.5112
+iaf_20231213_2 2.96 0.0696 2.97 0.0658
+iaf_20231213_3 4.96 0.0783 2.79 0.0537
+if_20231213_4 3.23 0.0845 2.48 0.1233
+if_20231213_5 3.57 0.1112 4.60 0.1564
+Avg. 3.72 0.1618 3.26 0.1821
+ARS548 $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
+st_20231213_1 3.02 0.4107 1.79 0.3561
+iaf_20231213_2 2.62 0.0492 3.37 0.0529
+iaf_20231213_3 4.07 0.0418 2.67 0.0380
+if_20231213_4 2.02 0.0805 2.37 0.0683
+if_20231213_5 1.93 0.0825 3.09 0.0902
+Avg. 2.73 0.1329 2.66 0.1211
+V-B Scan matching sensitivity study
+Afterwards, we individually test the stability and sensitivity of the Gaussian scan matching algorithm.
+For this purpose, we select the if_20231208_4 sequence from Snail-Radar due to its length (2.23 km over 8 min 35 s) and its challenging urban geometry.
+We extract radar scans (from both radars) every 30 s, which are then modeled with every algorithm (NDT, GICP, VGICP, Gaussians).
+We perform scan matching experiments between every model and 401 transformed copies of their source radar point clouds: 1 using the identity transformation, 100 using random translations, 100 random rotations, 100 random combined translations/rotations, and 100 copies of the scan with added noise.
+The random translations and rotations are uniformly sampled between 0 and 10 m/0°and 10°respectively.
+The direction of translation and axis of rotation is also uniformly sampled.
+The noise is normally distributed around zero with $\sigma=1$ m.
+Figure 2 shows results for both Oculii Eagle and ARS548 radar scans.
+We can observe that our approach produces much lower translation and rotation errors across the board, with a much more compact error distribution.
+Failure rate is also improved compared to other algorithms in the case of random translations, combined transforms or noise.
+While it is higher in random isolated rotations, we believe this contributes to rejecting badly converged solutions and thus it can be said that our algorithm behaves more robustly.
+We must also mention that even though the NDT implementation is capable of reporting convergence failure this never seems to happen (failure rate of 0%), and thus the errors are much higher and spread out than for the other methods.
+We report a single metric for our method in this experiment.
+This is because nearly all particles converge to the same solution, resulting in our algorithm obtaining the same metrics regardless of particle count.
+We think this happens because the same scan is used for both modeling and scan matching, and thus the likelihood of converging to a different local minimum is reduced.
+This also shows a strength of our algorithm: under “optimal” circumstances the aforementioned near full convergence of all particles occurs even though particle dispersion is high.
+The next experiment section is dedicated to a full odometry pipeline, where subsequent scans are registered against a model from a previous scan, which allows the multi-hypothesis strategy to take effect.
+V-C Odometry experiments
+Finally, we perform a comprehensive evaluation on a full Radar-Inertial Odometry task using both datasets and radar types.
+For this purpose we develop our own odometry pipeline following well established Extended Kalman Filter practice [31, 30, 12] to tightly integrate all incoming sensor information.
+Leveraging the Doppler information provided by 4D radar, scans are preprocessed with the RANSAC-LSQ algorithm [30] to both filter out outlier (dynamic) points, and estimate the current radar-frame velocity of the robot (egovelocity).
+The filter state includes position, velocity, attitude, accelerometer/gyroscope bias, and radar-to-IMU pose.
+We perform Kalman propagation using inertial data, and Kalman updates using radar egovelocity observations and scan matching results.
+In particular for the scan matching we follow [32] and constrain the observation model by only considering 3 degrees of freedom: X/Y/yaw.
+This is justified in the fact that radars have lower resolution in the Z axis, which degrades the usefulness of Z/roll/pitch information returned by scan matching algorithms.
 We evaluate relative translation/rotation errors using the modern and flexible evo evaluation package [33], while preserving existing practice by using a custom evaluation script that follows the behavior of older packages such as rpg_trajectory_evaluation as closely as possible.
 In the case of Snail-Radar sequences, we report rotation errors in mrad/m instead of º/m due to much smaller errors.
-Besides our Gaussian methods, we evaluate four ablations within our RIO framework: one without multi-hypothesis registration, and three using common established scan matching algorithms (NDT, GICP and VGICP) directly with radar point clouds.
-We use the current nominal state of the EKF directly as the initial estimate for single-hypothesis scan matching.
-The NDT version uses the implementation provided by the well known PCL (Point Cloud Library), while the GICP and VGICP versions use small_gicp [34] as the underlying implementation.
-In the case of Oculii Eagle scans, we set a target of 20 points per Gaussian; while for ARS548 scans, due to the much lower number of returned points, we set a target of 8 points per Gaussian.
-In both cases we sample a swarm of 8 particles during scan matching, which we believe are enough for testing purposes without significantly increasing processing times.
-Other parameters include $t_{\mathrm{max}}$ = 15m and $\alpha_{\mathrm{max}}$ = 5° for keyframing, $d_{\mathrm{max}}$ = 4 for scan matching, and 0.5m/2° for multi-hypothesis particle dispersion.
-These parameters were chosen to suit the typical depth range/sparsity relation of a radar.
 The source code of our system, which includes the Gaussian algorithms and the odometry pipeline, can be accessed online on GitHub [^1] https://github.com/robotics-upo/gaussian-rio-cpp .
@@
 The code also contains all details regarding the system parameters we used, such as EKF covariance initialization, IMU/noise parameters, etc.
-VI-A Results on NTU4DRadLM
-We evaluate using the NTU4DRadLM dataset [8], which includes scans captured by an Oculii Eagle radar, and a VectorNav VN-100 IMU.
+V-C1 Results on NTU4DRadLM
+We evaluate using the NTU4DRadLM dataset [8].
 Besides the evaluation performed with our pipeline, we also include previously published results of several radar-focused odometry methods, in order to provide context on the current performance of the state of the art.
@@
 Table I shows the comparison between all methods.
-Our Gaussian methods provide the best or second best results for all sequences except for loop3, which seems to favor GICP-based methods (both baseline and our implementation).
-Moreover, in 3 out of the 4 sequences the multi-hypothesis scan matching approach provides improvements in translation error, especially in cp.
+Our Gaussian method provides the best or second best results for all sequences except for loop3, which seems to favor GICP-based methods (both baseline and our implementation).
+Performing multi-hypothesis scan matching helps in some sequences such as cp and loop3.
+Moreover, a high number of particles is not necessary to obtain these improvements.
 Additionally, we report the average scan matching processing time for each method running under our odometry pipeline.
@@
 The single-hypothesis version of our method achieves the fastest scan matching processing times, closely followed by VGICP.
-The multi-hypothesis approach, despite simultaneously registering 8 particles at once, is only 3.62x slower than the single-hypothesis version.
+When registering multiple particles our method shows non-linear scaling, as 8 particles only need 3.62x more time than a single particle, while 128 particles need 82x more time.
 We also mention that our current RIO implementation is entirely sequential, and that Gaussian modeling at keyframes consumes up to 608.78ms, raising the average scan processing time to 18.13ms.
 However, we believe that running the modeling and odometry tasks in parallel would remove any timing inconsistencies affecting real-time usage; which we leave as future work.
-VI-B Results on Snail-Radar
-In addition to NTU4DRadLM, we perform additional testing on the Snail-Radar dataset [32], which uses both Oculii Eagle and ARS548 radars together with an Xsens MTi-3-DK IMU, and includes sequences recorded in harsher environment conditions.
+V-C2 Results on Snail-Radar
+In addition to NTU4DRadLM, we perform additional testing on the Snail-Radar dataset [28].
 In particular, we selected the five sequences recorded on 2023/12/13 because of the nighttime and rain conditions of that day under a high speed car platform.
@@
 In general, we can observe the best results to be spread across the different methods.
-Nonetheless, our Gaussian algorithms obtain the best or second best results in 9 out of 10 metrics, and the best translation error in 3 out of 5 sequences (the other two favoring NDT).
-The multi-hypothesis approach improves or matches results compared to the single hypothesis version in 3 out of 5 sequences.
+Nonetheless, our Gaussian method obtains the best or second best results in 8 out of 10 metrics, and the best translation error in 3 out of 5 sequences (the other two favoring NDT).
+Our Gaussian approach produces the best translation errors in 3 out of 5 sequences, and the best rotation errors in 2 out of 5 sequences.
+The multi-hypothesis approach is capable of improving performance in 4 out of 5 sequences.
+Regarding the number of particles, a higher number benefits 3 out of 5 sequences, which indicates much more challenging geometry compared to NTU4DRadLM.
 We also note that NDT seems to behave much better in Snail-Radar than in NTU4DRadLM.
@@
 However, we can observe better metrics in general compared to Oculii Eagle.
-The sparser scans trigger a known issue in PCL’s implementation of NDT that cause a program termination, so unfortunately NDT is excluded from the comparison.
-The results once again show different methods leading in different sequences, however it is worth pointing out that 6 out of the 10 best results are now obtained by our Gaussian algorithms, as well as 3 out of the 10 second best results.
-Another thing to mention is that the multi-hypothesis scan matching seems to perform worse than the single-hypothesis version.
-We hypothesize this to be related to the increased sparsity of the scans, and believe it could be solved by improving the sampling distribution of the particles.
-VI-C Additional qualitative experiments
-In order to further justify our modeling approach based on globally optimized 3D Gaussians versus existing voxel-based approaches, we compare in Fig.
-2 the output of our method against NDT.
-To do this, we prepare a radar-based map of NTU4DRadLM’s cp sequence using the supplied ground truth, and model a section of said map with both NDT and our method.
-The positions and sizes of the Gaussians generated by NDT are clearly constrained and influenced by the voxel grid, while our approach generates better fitted Gaussians.
-Moreover, our method is able to preserve geometric details such as corners, and allocates a higher budget of Gaussians to areas with higher point density, resulting in overall better representation of the geometry.
-Additionally, we include a video supplement.
-In the video we show several Gaussian models of the same map created using different points-per-Gaussian ratios.
-It can be seen that, regardless of said hyperparameter, our modeling method is able to generate a representation reasonably fitting the map.
-Different number of Gaussians are allocated to areas of the map depending on the density and complexity of the area, as well as the available budget.
-Moreover, when the ratio is higher (meaning fewer Gaussians), sparse regions with few (outlier) radar points do not receive any Gaussians.
-Finally, we also include an example of the radar-inertial odometry pipeline on the same NTU4DRadLM cp sequence, compared to the ground truth trajectory.
-Figure 2: Qualitative comparison between NDT and our Gaussian modeling using NTU4DRadLM’s cp sequence.
-VI-D Summary of results
+The results once again show different methods leading in different sequences, however it is worth pointing out that our method produces the best results in 2 out of 5 sequences.
+V-C3 Ablation study
+In order to show the individual contribution of our global, jointly optimized Gaussian modeling approach, we perform an ablation study based on the odometry experiment combining our scan matching implementation with a strong locally constructed/voxelized baseline for modeling (VGICP).
+That is, we use the Gaussian models generated by VGICP with our scan matching approach and a single particle.
+Table III shows the results on all tested sequences.
+In the case of NTU4DRadLM our global models clearly obtain the best metrics on all sequences.
+For Snail-Radar the improvements are more marginal and reduced to some sequences, however our approach still obtains the best translation errors on average with both radars, while obtaining the best rotation errors on average with ARS548.
+V-C4 Qualitative results
+Fig.
+3 shows plots of select sequences from both datasets.
+Specifically, we plot the loop2 sequence from NTU4DRadLM and the iaf_20231213_3 sequence from Snail-Radar, the latter for both Oculii Eagle and ARS548 radars.
+These two sequences are especially long and complex, which allows for qualitatively evaluating the performance of the different methods.
+We can observe that NDT drifts the most in loop2, while the rest stay fairly close to each other, and in particular our algorithm with 128 multi-hypothesis particles is the closest to the ground truth, closely followed by the 8 particle and 1 particle (single hypothesis) versions.
+As for iaf_20231213_3, the algorithms generally perform better with the ARS548.
+The 128 particle version of our algorithm is especially close to the ground truth using both radars.
+V-D Summary of results
 We consider that the results above validate our methodology.
-Our Gaussian modeling and scan matching algorithm produces comparable relative translation and rotation metrics to existing registration algorithms, as shown by the experiments.
+Our Gaussian modeling and scan matching algorithm produces comparable or superior relative translation and rotation metrics to existing registration algorithms, as shown by the experiments.
 Moreover, in several sequences across several datasets and radar types our algorithm produces the best overall results.
-On the other hand, the multi-hypothesis approach seems to need further adjustments, although it is already capable of producing better results for some sequences.
+On the other hand, although the multi-hypothesis approach provides improvements in some sequences, it also produces regressions in others, indicating the need for further adjustments.
 As an additional observation, it is worth mentioning that we believe the radar-to-IMU calibration file provided by NTU4DRadLM to be inaccurate, as it does not seem to match the placement of the different sensors in the photo they provide of their platform.
 The effects of this inaccurate calibration can be especially seen in the rotational errors, which are much higher for all methods compared to those achieved under Snail-Radar sequences.
-VII Conclusion
-In this paper we introduce modeling and scan matching algorithms based on 3D Gaussians, which initialize and jointly optimize all parameters of a 3D Gaussian model to fit a radar point cloud, and make use of multiple hypothesis to support a scan matching process for subsequent incoming point clouds.
-Through experiments in multiple datasets with multiple radar types, we show that these methods produce results comparable to existing registration algorithms, and even manage to outperform them in several sequences.
-At the same time, our methods offer higher flexibility/expandability and faster scan matching compared to existing approaches based on individual points and voxelization [2, 1, 15, 16].
-Future works include improving the run-time efficiency of the Gaussian modeling (replacing gradient descent, and running in parallel to the odometry pipeline), better initialization and dynamic management of Gaussians (as in 3DGS), as well as incorporating radar-specific information such as RCS (Radar Cross Section) values, and improving the robustness of the multi-hypothesis particle system with kinematic priors.
-Another line of work is further development of the RIO pipeline, using techniques such as Pose Graph Optimization (PGO), loop closure, and creating a global Gaussian map of the environment for the purposes of full SLAM.
-Acknowledgment
-This work was partially supported by the following grants: 1) INSERTION PID2021-127648OB-C31, and 2) NORDIC TED2021-132476B-I00 projects, funded by MCIN/AEI/ 10.13039/501100011033 and the “European Union NextGenerationEU / PRTR”.
+VI Conclusion
+In this paper we introduce a new modeling approach for radar point clouds based on globally and jointly optimized 3D Gaussians unlike existing voxelized or locally constructed approaches [1, 2, 15].
+We show how this representation produces more robust and better performing scan matching, which we also propose to perform using multiple registration hypotheses.
+Through odometry experiments in multiple datasets with multiple radar types, we show that our approach produces results comparable to existing registration algorithms, managing to outperform them in several sequences.
+At the same time, our method offers higher expandability and faster scan matching compared to existing locally constructed or voxelized approaches [2, 1, 15, 16].
+Future works include improving the run-time efficiency of the Gaussian modeling (replacing gradient descent, and running in parallel to the odometry pipeline), dynamic management of Gaussians (as in 3DGS), as well as incorporating radar-specific information such as RCS (Radar Cross Section) values, and improving the formulation of the multi-hypothesis particle system with kinematic priors.